### v2

# ProReason: Multi-Modal Proactive Reasoning with Decoupled Eyesight and Wisdom

Abstract
Large vision-language models (LVLMs) have witnessed significant progress on visual understanding tasks. However, they often prioritize language knowledge over image information on visual reasoning tasks, incurring performance degradation. To tackle this issue, we first identify the drawbacks of existing solutions (i.e., limited multi-modal reasoning capacities, and insufficient and irrelevant visual descriptions). We then decompose visual reasoning process into two stages: proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), and introduce a novel visual reasoning framework named ProReason. This framework features decoupled vision-reasoning capabilities and multi-run proactive perception. Briefly, given a multi-modal question, ProReason iterates proactive information collection and reasoning until the answer can be concluded with necessary and sufficient visual descriptions. Notably, the disassociation of capabilities allows seamless integration of existing large language models (LLMs) to compensate for the reasoning deficits of LVLMs. Our extensive experiments demonstrate that ProReason outperforms both existing multi-step reasoning frameworks and passive peer methods on various benchmarks for both open-source and closed-source models. More impressively, with the assistance of LLMs, ProReason achieves a performance improvement of up to 15% on MMMU benchmark. Our insights into existing solutions and the decoupled perspective for feasible integration of LLMs illuminate future research on visual reasoning techniques, especially LLM-assisted ones [^1] Code is available at https://github.com/lian-tian-mo-zun/Pro_Reason.
ProReason: Multi-Modal Proactive Reasoning
with Decoupled Eyesight and Wisdom
Jingqi Zhou, Sheng Wang, Jingwei Dong, Lei Li, Jiahui Gao, Jiyue Jiang, Lingpeng Kong, Chuan Wu The University of Hong Kong, The Chinese University of Hong Kong, Xi’an Jiaotong University {u3011211, u3009618}@connect.hku.hk, dongjingwei@stu.xjtu.edu.cn, nlp.lilei@gmail.com, sumiler@connect.hku.hk, jiangjy@link.cuhk.edu.hk, {lpk, cwu}@cs.hku.hk
1 Introduction
In recent years, large language models (LLMs) (Dubey et al., 2024; Team et al., 2023; Jiang et al., 2023) have experienced explosive growth in their capabilities, driving significant advancements across various fields (Shao et al., 2023; Guo et al., 2024; Shao et al., 2024). This progress has also sparked interest in developing large vision-language models (LVLMs) (Chen et al., 2024a; Bai et al., 2023), which, like LLaVA (Li et al., 2024b), have achieved remarkable performance in multi-modal understanding tasks. However, state-of-the-art (SOTA) LVLMs still struggle to handle visual understanding with textual reasoning simultaneously due to inherent modality differences. For example, Ghosh et al. (2024) demonstrate that LVLMs often rely more on their prior language knowledge, neglecting visual information in multi-modal reasoning tasks such as visual chart understanding and math reasoning, resulting in performance degradation. Figure 2.b illustrates a typical case of this issue, where the reasoning process remains irrelevant to the image.
To address the challenges, a promising solution is to extract visual information from images into textual form to assist LVLMs in reasoning (Yin et al., 2023b; Mitra et al., 2024). For example, Ghosh et al. (2024) instruct LVLMs to generate fine-grained captions to facilitate the subsequent reasoning process. However, these existing methods exhibit two primary limitations: (i) The visual extraction process is question-agnostic and reasoning-free, that the image description is not targeted for a given question, and reasoning is not involved to deduce extra information for better descriptions. This drawback, termed as “passive”, results in irrelevant or insufficient information, and subsequent performance degradation. (ii) The reasoning process entangles visual understanding with textual reasoning abilities of a single LVLM. As mentioned above, the limited capabilities of LVLMs do not always successfully handle both abilities and produce high-quality reasoning. Hence, both of these limitations hinder the performance of LVLMs in multi-modal reasoning tasks.
To address both shortcomings, we propose a novel multi-step multi-modal reasoning framework named ProReason, which features disentangled vision-reasoning capabilities and proactive (i.e., question-oriented and reasoning-involved) visual extraction. As illustrated in Figure 1, we first decouple multi-modal reasoning capacity into two sub-tasks: proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), based on the capacity of LVLMs. The former focuses on understanding images and converting visual details into textual form, while the latter integrates the gathered information to draw final conclusions, with each involving different agents. In the proactive visual perception stage, a Dispatcher decomposes the original question into visual or reasoning-related sub-questions and selectively engages a Vision Expert to capture additional visual information, or a Reasoning Expert to derive extra insights. Subsequently, a Referee evaluates the sufficiency of known information, deciding whether to continue the proactive visual perception stage or proceed to the textual reasoning stage, where all information is consolidated for a Summarizer to generate the final answer. Unlike passive methods, all sub-agents operate based on the given question and current descriptions, effectively avoiding information redundancy or insufficiency. Notably, the disentangled vision-reasoning does not require vision-irrelevant roles (i.e., Reasoning Expert, Referee, and Summarizer) to be performed by LVLMs, enabling the seamless integration of existing LLMs with proven strong reasoning abilities (Chang et al., 2024), thereby alleviate the limitations of LVLMs.
Empirically, we evaluate ProReason across four challenging visual reasoning benchmarks with both open-source and closed-source models. Extensive experiments demonstrate that ProReason achieves significant and consistent performance improvement over both existing reasoning frameworks and state-of-the-art passive methods across all the benchmarks, with a peak enhancement as 13.2% (Sec. 4.2). Besides, our comparative analysis highlights the necessity and relative importance of each sub-agent (Sec. 4.3), as well as the crucial role of decoupling (Sec. 4.4). With the verified superiority over the simultaneously inherent usage of both capabilities, the decoupled perspective not only demonstrates that textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks (Sec. A.3.1), but also exemplifies the feasibility of LLM-assisted LVLM reasoning, exhibiting performance improvements of up to 15% on the MMMU benchmarks (Yue et al., 2023).
The main contributions of this work are threefold:
  1. We propose a novel multi-modal reasoning framework named ProReason, featuring decoupled vision-reasoning and iterative proactive perception capabilities.
  2. Empirically, we identify the drawbacks of existing solutions (i.e., limited multi-modal reasoning capacities, and insufficient and irrelevant visual descriptions), and validate the successful mitigation of these issues with ProReason by the superior performance over peer methods.
  3. ProReason showcases the remarkable feasibility of integrating LLMs for boosted visual reasoning, illuminating the potential for LLM-assisted LVLM reasoning in future research.
2 Preliminary Observations
We first analyze the behaviors of several existing methods on the MMMU (Yue et al., 2023) dataset, a challenging multi-modal benchmark requiring comprehensive college-level knowledge and fine-grained reasoning abilities. All experiments are conducted with three recent LVLMs for the robustness of the results: Llama3-LLaVA-NeXT-8B (Li et al., 2024a), LLaVA-OneVision-Qwen2-7B-OV (Li et al., 2024b) and Qwen-VL-Chat (Bai et al., 2023). Further details are provided in Sec. 4.1.
2.1 Challenges in Visual Reasoning: Limitations of LVLMs
Chain-of-Thought(CoT) (Wei et al., 2022) has been extensively verified to enhance the performance of LLMs. Here, we explore its impact on the reasoning performance of LVLMs. Similarly, we require the models to “think step by step” before generating final answers. As shown in Table 8, a counter-intuitive phenomenon is observed: compared to “Direct” answering method, the introduction of CoT consistently incurs slight performance degradation across all three models. Inspired by the findings of Ghosh et al. (2024) that LVLMs often rely on prior language knowledge yet neglecting visual information, we hypothesize that the degraded performance might be caused by the image-irrelevant reasoning of CoT.
To validate the hypothesis, we design a pipeline inspired by Liu et al. (2023c) to measure the correlation between model responses and image contents. Specifically, given a question, we take the reasoning process generated by GPT-4o-mini [^2] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence, whose superiority on MMMU task is verified in Sec. 4.2, as the golden solution. Subsequently, a more powerful judge (i.e., GPT-4 (Achiam et al., 2023)) is employed to evaluate the relevance of responses of candidate models (e.g., Llama3-LLaVA-NeXT-8B) and this golden solution with the prompt template in Figure 8, and gives a comprehensive score from 1 to 5, with 5 indicating the highest relevance. Since the golden solution contains the essential information needed to solve the problem within the image, a higher correlation indicates a stronger relevance to the image content. As shown in Table 9, the reasoning chains of incorrect responses are significantly less relevant to the images than those of correct answers. An exemplary case is shown in Figure 2, where Llama3-LLaVA-NeXT-8B attempts to use the right-hand rule to solve the task, but its reasoning process is almost unrelated to the image, ultimately leading to an incorrect conclusion. This highlights the limitation of CoT for multi-modal reasoning tasks, that image-irrelevant reasoning process provokes performance degradation, since the limited capabilities of LVLMs cannot effectively handle both image and text information simultaneously.
2.2 Drawbacks of Passive Information Extraction
Passive visual reasoning techniques suffer insufficient and irrelevant visual information, despite mitigating the oversight of images by converting them into detailed captions, like Visual Description Grounded Decoding (VDGD) (Ghosh et al., 2024). To support this claim, we generate fine-grained image captions using GPT-4o-mini with the prompt shown in Figure 9. We then incorporate these captions into the prompts for LVLMs to facilitate the reasoning process. As listed in Table 8, while these image descriptions improve the performance of LVLMs, the gains are marginal, consistently amounting to less than 1%. This underscores the limited utility of captions generated by passive methods.
For further demonstration, we analyze the generated captions along three dimensions: Detail Level, Question Relevance, and Reasoning Effective Info Inclusion, measuring the richness of detail, relevance to the given question, and the inclusion of information that is essential for reasoning, respectively. Similar to Sec. 2.1, we pair the image captions with the reasoning process of GPT-4o-mini, and employ GPT-4 as a judge to rate the captions from 1 to 5 across these dimensions. The prompt template is shown in Figure 8. As shown in Table 13, the captions for correct responses of Llama3-LLaVA-NeXT-8B receive higher scores across all three criteria, highlighting the importance of better captions for multi-modal reasoning. Additionally, all captions score significantly lower in the Question Relevance and Reasoning Effective Info Inclusion dimensions than the Detail Level dimension, indicating that while the captions are detailed, they often lack relevance to the questions. Figure 2 shows a case where Llama3-LLaVA-NeXT-8B utilizes fine-grained captions to solve a question from the MMMU benchmark. As illustrated, although the caption exhaustively describes the image content, it incorrectly describes the wires in the image as octagons, and misses information about the locations of these wires. This information is irrelevant to the target question, thus offering minimal assistance to LVLMs. In summary, our analysis highlights the drawbacks of passive visual reasoning enhancement techniques in terms of information insufficiency and redundancy, due to their question-agnostic property.
Figure 1: Overview and comparison of ProReason, VDGD and ReAct. Unlike existing works (e.g., VDGD and ReAct), our proposed method decouples visual perception and textual reasoning while allowing the model to actively acquire necessary information from the images, achieving superior performance.
(1) ProReason
(2) Passive Visual Reasoning Enhancement Technique
(3) Multi-step Reasoning Framework for LLMs
3 Method
As depicted in Figure 1, ProReason presents an innovative decoupling of the visual reasoning process into two distinct phases framed through the lens of LVLM capabilities: Proactive Visual Perception (i.e., eyesight) and Textual Reasoning (i.e., wisdom). The entire workflow consists of five functionally distinct yet inter-cooperative sub-agents, along with a Memory component, facilitating multi-modal reasoning performance.
3.1 Proactive Visual Perception
Proactive Visual Perception is the core of question-oriented visual information extraction, driven by four sub-agents: Dispatcher, Vision Expert, Reasoning Expert and Referee. The Dispatcher breaks down the original question, selectively directing the Vision Expert to capture specific visual information, or instructing the Reasoning Expert to analyze known information to derive more. This results in the decoupling of capabilities within the LVLMs, with the Vision Expert handling visual perception and the Reasoning Expert addressing textual reasoning. The responses from both experts are stored in a textual Memory component. The Referee then evaluates whether the information stored in Memory is sufficient to answer the original question.
Formally, given an image $I$ and its corresponding textual question $Q$, the Dispatcher decides to consult the Vision Expert or Reasoning Expert, based on the analysis of $Q$ and the known information in the Memory (if not empty). The Dispatcher then generates a query $q$ for the chosen expert. If the Vision Expert is selected, it takes the image $I$ and query $q$ as input, and generates an answer $A_{v}$, which is then stored in the Memory. When the Reasoning Expert is selected, it provides a response consisting of the reasoning process and the final answer $A_{r}$ based on the query $q$ and known information in the Memory, before only $A_{r}$ is stored in Memory. The Referee then evaluates the available information in the Memory concerning the question $Q$. If the Memory contains adequate information to answer the question $Q$, the Referee outputs the identifier “SOLVABLE”; otherwise, it outputs “UNSOLVABLE”. If the Referee’s output is SOLVABLE, the workflow precedes to the Textual Reasoning phase. Conversely, if the output is UNSOLVABLE, the above process will be re-executed to gather more necessary information.
In the Proactive Visual Perception phase, four sub-agents collaborate closely to breakdown visual reasoning tasks, thereby achieving an initial decoupling of LVLMs’ visual and reasoning capabilities. Simultaneously, this approach enables the framework to proactively acquire the necessary information and prevent omissions, thereby overcoming the drawbacks of passive methods. Notably, the Memory component allows ProReason to keep compact information, and avoids lengthy reasoning traces like CoT (Wei et al., 2022) and ReAct (Yao et al., 2022), thereby suffering less from redundant information (for a detailed analysis, refer to Sec. A.3.1 and Table 13).
3.2 Textual Reasoning Step
The Textual Reasoning step focuses on integrating the available information in the Memory, and providing the final answer to the question $Q$. This step is mainly powered by a sub-agent called Summarizer. Once the Referee determines that the information in the Memory is sufficient to address the question $Q$, and outputs the identifier SOLVABLE, the Summarizer will be called to draw a conclusion, based on the Memory information. This conclusion represents the answer to the question $Q$ of the whole framework, and will be evaluated by the performance metrics.
3.3 Advantages of ProReason
Decoupled Visual Reasoning.
In ProReason, the complete multi-modal reasoning process is decomposed into visual perception and textual reasoning stages, each executed by separate agents. These agents are then effectively organized through a designated pipeline, significantly enhancing the ability of LVLMs to tackle visual reasoning tasks.
LLM-assisted multi-modal reasoning.
On the other hand, this decomposition-before-integration approach not only circumvents the inherent differences between modalities but also allows visually irrelevant sub-agents to be performed by text-only LLMs, thereby facilitating LLM-assisted multi-modal reasoning. Consequently, the extensively verified powerful reasoning capabilities of LLMs can be seamlessly synergized to power the multi-modal reasoning process for performance enhancement.
Reduced information Mission or Redundancy.
With the close collaboration of the sub-agents in Proactive Visual Perception step, ProReason can proactively (i.e., question-orientedly and reasoning-involvedly) extract the necessary and sufficient visual details from images, effectively avoiding information omission or redundancy. Meanwhile, the Memory component retains only the image captions from the Vision Expert and the reasoning results from the Reasoning Expert, providing a compact textual descriptions. This contributes to minimize the interference of irrelevant information on the subsequent Summarizer.
4 Experiments
In this section, we first evaluate the performance of our ProReason framework against recent baselines on multiple benchmarks, followed by an in-depth ablation analysis of different components.
4.1 General Setup
Datasets.
To comprehensively validate the performance of our framework, we conduct experiments across four benchmarks: Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023a)  [^3] Due to our emphasis on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning. To facilitate the comparison across different benchmarks, the results for MME benchmark are calculated by the percentage of correct answers out of the total answers. For the results using the official MME calculation method, please refer to Table 16, Table 16, Table 18, Table 16 and Table 18. , Massive Multi-discipline Multi-modal Understanding and Reasoning (MMMU) (Yue et al., 2023), MathVista (Wang et al., 2024), and HallusionBench (Liu et al., 2023a). All of them require visual reasoning capabilities to complete the tasks correctly, and are introduced briefly in Sec. A.2.
Base Models.
We employ GPT-4o-mini, Llama3-LLaVA-NeXT-8B [^4] The model is deployed using CUDA https://developer.nvidia.com/cuda-toolkit on a NVIDIA A100 graphics cards. and Qwen2.5-72B- Instruct (Team, 2024) as our base models, due to their excellent performance, accessibility, and representativeness of model sources. As one of the most performant LVLMs, GPT-4o-mini demonstrates significant advancements in visual reasoning capabilities, and provides cheap and fast API. In contrast, Llama3-LLaVA-NeXT-8B is a fully open-sourced LVLM developed by the LLaVA team based on the Llama-3-8B LLM (Dubey et al., 2024) and CLIP vision encoder (Radford et al., 2021). Qwen2.5-72B-Instruct is a robust LLM designed to deliver high-quality instruction following and handle complex tasks. It will be utilized for achieving LLM-assisted multi-modal reasoning.
Baselines.
Besides the most basic method where models are instructed to answer questions directly, we compare ProReason with two categories of peer methods. To determine the benefits of proactive information extraction in ProReason, we first consider two SOTA passive visual reasoning methods, VDGD (Ghosh et al., 2024) and CCoT (Mitra et al., 2024). Additionally, we choose two multi-step reasoning frameworks of LLMs, Chain of Thought (CoT) (Wei et al., 2022) and ReAct (Yao et al., 2022), to demonstrate the effects of directly migrating LLM solutions to LVLMs.
  1. Direct. As indicated by the name, models are required to answer questions directly without dedicated prompts. This baseline is set to evaluate the initial performance of base models.
  2. VDGD. With image caption prefixed to text instruction, VDGD selects tokens that deviate the least from the description using a formula based on Kullback-Leibler divergence.
  3. CCoT. Given an image and the question, CCOT first generates a scene graph of the image with LVLMs, and then extracts the answer by prompting the LVLMs with the graph.
  4. CoT. CoT is an advanced prompting method that encourages LLMs to break complex tasks down into a series of easy steps, which has been applied broadly and verified to boost the reasoning performance remarkably (Chu et al., 2023).
  5. ReAct. ReAct is an LLM-specific agent framework, which performs tasks by alternating between reasoning and execution behaviors. To extend it to multi-modal domain, we use two LVLMs to perform both steps, and rename them as the Vision and Reasoning Experts, respectively. This aligns with our notions for easy understanding, and is shown in Figure 1.
Implementation Details.
The prompt templates for all methods are shown in Figures 9, 10 and 12. Specifically for ProReason, to prevent infinite loops, if the Dispatcher selects the Vision Expert or Reasoning Expert to obtain information up to 5 consecutive times, and the Referee still determines that the existing knowledge in Memory remains insufficient to resolve the question, the Memory will be cleared to restart the information acquisition process. If the system fail to break the loop after 5 attempts, the proactive visual perception phase will be immediately terminated. Subsequently, the Summarizer will generate the final answer based on the available information in Memory from the last attempted iteration. This setup, refined through multiple trials, is the most effective. Additionally, since we cannot obtain the tokens output by GPT-4o-mini, we omit the step of selecting the token with the smallest deviation from the image description when implementing VDGD for GPT-4o-mini.
Table 1: Performance of multiple approaches with two base models across four visual reasoning benchmarks. “Hallu.” is the abbreviation of HallusionBench. Based on the performance of the direct method, red and blue signify the improvement and degradation, respectively.
Model Method Dataset
MME MMMU MathVista Hallu.
Llama3-LLaVA- NeXT-8B Direct 61.5 41.8 37.1 45.8
VDGD 68.8 (+7.3) 42.3 (+0.5) 36.1 (-1.0) 44.2 (-1.6)
CCoT 68.9 (+7.4) 40.5 (-1.3) 36.8 (-0.3) 37.4 (-8.4)
CoT 58.8 (-2.7) 41.5 (-0.3) 35.9 (-1.2) 43.1 (-2.7)
ReAct 68.5 (+7.0) 46.7 (+4.9) 31.7 (-5.4) 43.6 (-2.2)
ProReason 71.5 (+10.0) 50.5 (+8.7) 38.8 (+1.7) 50.9 (+5.1)
Average 66.30 43.9 36.06 44.16
GPT-4o-mini Direct 79.2 48.4 53.0 56.0
VDGD 82.3 (+3.1) 51.4 (+3.0) 51.2 (-1.8) 52.4 (-3.6)
CCoT 80.8 (+1.6) 54.2 (+5.8) 53.6 (+0.6) 56.7 (+0.7)
CoT 87.8 (+8.6) 58.5 (+10.1) 53.8 (+0.8) 56.3 (+0.3)
ReAct 87.3 (+8.1) 54.8 (+6.4) 49.3 (-3.7) 51.1 (-4.9)
ProReason 91.9 (+12.7) 61.6 (+13.2) 54.9 (+1.9) 59.9 (+3.9)
Average 84.63 54.82 52.63 55.4
Table 2: Performance of Llama3-LLaVA-NeXT-8B Vision Expert, assisted by different text-only LLMs, on various benchmarks under ReAct and ProReason frameworks. The red and blue texts indicate the improvements and reductions relative to the case of using the same method without the assistance of text-only LLMs, respectively.
Vision Expert Textual Sub-Agents Method Dataset
MME MMMU MathVista Hallu.
Llama3- LLaVA- NeXT-8B Llama3-LLaVA- NeXT-8B Direct 61.5 41.8 37.1 45.8
ReAct 68.5 46.7 31.7 43.6
ProReason 71.5 50.5 38.8 50.9
Qwen2.5-72B- Instruct ReAct 71.0 (+2.5) 50.4 (+3.7) 34.6 (+2.9) 40.4 (-3.2)
ProReason 81.3 (+9.8) 56.8 (+6.3) 48.8 (+10.0) 52.3 (+1.4)
GPT-4o-mini ReAct 73.6 (+5.1) 48.4 (+1.7) 36.2 (+4.5) 46.7 (+3.1)
ProReason 84.7 (+13.2) 54.5 (+4.0) 41.7 (+2.9) 53.1 (+2.2)
4.2 Main Results
ProReason exhibits significant and consistent performance enhancement over baselines across all the benchmarks. As listed in Table 1, despite better performance than the direct method on MME dataset, VDGD and CCoT fail to demonstrate consistent improvements on the other datasets. In contrast, ProReason consistently outperforms all the other baselines for both Llama3-LLaVA-NeXT-8B and GPT-4o-mini model across all benchmarks, with a peak improvement as 13.2%, demonstrating the superiority and task robustness of ProReason.
Proactive information acquisition surpasses SOTA passive methods, especially in complex visual reasoning tasks. Specifically, compared to MME, MathVista and HallusionBench present higher image complexity and question difficulty, and thus require stronger visual understanding and textual reasoning capabilities. This leads to performance degradation of passive methods (i.e., VDGD and CCoT), highlighting their limited applicability to complex visual reasoning tasks. In contrast, ProReason achieves notable performance improvements, up to 5.1%, by proactively acquiring visual information from images rather than generating question-agnostic captions. This aligns with out previous obervations in Sec. 2.2 that passive methods introduce substantial information redundancy or omission, misleading subsequent reasoning processes.
Decoupling the visual perception and textual reasoning capabilities of an LVLM outperforms their simultaneous inherent usage. As listed in Table 1, when both capabilities are utilized concurrently, CoT consistently degrades performance compared to the “Direct” method across all benchmarks with Llama3-LLaVA-NeXT-8B model, consistent with the findings in Sec. 2.1. In contrast, despite the same models, ProReason alternates between visual information acquisition and textual reasoning processes, allowing to leverage each capability more effectively. This enables ProReason to consistently outperform CoT with both Llama3-LLaVA-NeXT-8B and GPT-4o-mini across all benchmarks, demonstrating the effectiveness of capability decoupling.
ProReason outperforms ReAct with even less token consumption. Given the similarity in multi-step reasoning, we compare ProReason and ReAct in terms of performance and cost. Specifically, as an LLM-specific multi-step reasoning framework, ReAct only outperforms the “Direct” method on MME and MMMU, but underperforms on MathVista and HallusionBench, showing inferior performance compared to the consistent improvements of ProReason. Furthermore, we compare their average token consumption of GPT-4o-mini. As shown in Table 13 of Appendix A, ProReason consumes significantly fewer tokens than ReAct on both MME and MathVista tasks, indicating its higher token efficiency and the importance of a compact Memory in reducing token usage. Coupled with better performance, this suggests the superiority of ProReason over LLM-specific ReAct framework.
Text-only LLMs can be effectively integrated into ProReason for dramatically enhanced performance. As mentioned in Sec. 3.3, the decoupled visual perception and textual reasoning capabilities facilitate the seamless integration of text-only LLMs. To demonstrate the utility of this advantage, we fix the Vision Expert as Llama3-LLaVA-NeXT-8B, and replace other agents with text-only LLMs. As listed in Table 2, with the assistance of either Qwen2.5-72B-Instruct or GPT-4o-mini, the Llama3-LLaVA-NeXT-8B Vision Expert receives remarkable performance boost across all benchmarks, particularly by 15% on MMMU and 11.7% on MathVista, compared to directly providing answers. In contrast, ReAct gains a much smaller improvement. This highlights the unique advantage of ProReason in leveraging existing text-only LLMs for enhanced performance. Notably, this advantage may open new avenues for continuously pushing the performance limits of LVLMs with the assistance of existing powerful LLMs.
4.3 Relative Importance of Sub-agents
To assess the importance of each sub-agent within the ProReason framework for visual reasoning tasks, we design five scenarios where Llama3-LLaVA-NeXT-8B acts as Dispatcher, Vision Expert, Reasoning Expert, Referee, or Summarizer, respectively, while the other sub-agents are powered by GPT-4o-mini. Given that Llama3-LaVA-NeXT-8B exhibits weaker visual understanding and textual reasoning capabilities than GPT-4o-mini, the more significant the performance drop incurred by replacing a sub-agent with Llama3-LaVA-NeXT-8B is, the more important that sub-agent is. Here we primarily consider the MME and MMMU benchmarks due to their comprehensive question coverage. The experimental results are presented in Table 13.
Summarizer is the most crucial sub-agent, closely followed by Referee. The replacement of Summarizer results in the most notable performance decline on both MME and MMMU tasks, reaching 6.2% and 10.6%, respectively. This highlights the critical function of the Summarizer in integrating all available information to conclude final answers. Besides, the substitution of Referee leads to a 10.1% reduction on MMMU. Given that MMMU is more challenging than MME, this finding underscores the essential role of the Referee in assessing the sufficiency of information, particularly in more complex visual reasoning tasks. The analysis in Section A.3.2 also demonstrates that the Referee plays a crucial role in enabling ProReason to accurately interpret visual detail information.
Relatively, Dispatcher and Reasoning Expert are the least essential sub-agents. Specifically, despite a decline, these two sub-agents exhibit significantly less performance degradation than other sub-agents. This can be attributed to the easier task of the Dispatcher, which requires minimal textual reasoning capabilities, and the infrequent calls of the Reasoning Expert, which is only activated when additional information needs to be inferred—a situation that is rare in current benchmarks. Besides, both sub-agents operate within the acquisition loop, allowing for greater error tolerance. Even if some error occurs, subsequent iterations can compensate for the missing information.
In summary, each sub-agent contributes to the performance of ProReason, underscoring their necessity. Relatively, the Summarizer and Referee are the most critical sub-agents, while the Dispatcher and Reasoning Expert have the least impact.
4.4 The Critical Implication of Decoupling
To validate the critical implication of decoupling visual perception and textual reasoning in ProReason while eliminating potential confounding factors from prompt engineering and multiple CoT implementations, we systematically integrate sub-agents through three configurations:
  1. Merge the Vision Expert and Reasoning Expert into a single sub-agent to examine the necessity of modality decoupling during sub-task execution.
  2. Integrate the Dispatcher, Vision Expert, Reasoning Expert, and Referee as a unified agent to verify the essentiality of the process design for the Proactive Visual Perception stage.
  3. Fully consolidate all five original sub-agents to demonstrate the pivotal role of decomposing visual reasoning tasks into distinct Proactive Visual Perception and Textual Reasoning phases.
The merged agents preserve identical prompts and maintain the same task execution procedures as their original counterparts, with their prompts shown in Figures 13, 14, and 15. Therefore, the performance degradation caused by agent merging quantitatively demonstrates the critical implication of decoupled processing in enhancing ProReason’s capabilities.
Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks. As shown in Table 13, merging Vision and Reasoning Experts results in a 3% performance drop on the MMMU benchmark, while combining Proactive Visual Perception with Textual Reasoning leads to a more significant 4.8% reduction (56.8 vs. CoT’s 58.5). Although agent merging also causes performance declines on MME, the merged versions still outperform CoT. Notably, given MMMU’s substantially higher complexity compared to MME, these findings reveal that the decoupling of visual-textual processing fundamentally drives its performance gains in complex scenarios.
5 Conclusion
In this paper, we first validate that existing multi-modal reasoning approaches still suffer insufficient and irrelevant visual descriptions, as well as limited multi-modal capacities. To address these issues, we decompose the visual reasoning process into visual perception and textual reasoning stages, and introduce a novel visual reasoning framework named ProReason, featuring decoupled vision-reasoning capabilities and multi-run proactive perception. Empirically, extensive experiments demonstrate the superiority of ProReason over both passive image information acquisition methods and multi-step reasoning frameworks for text-only LLMs across multiple visual reasoning benchmarks with both open-source and closed-source models. Notably, our method showcases the remarkable feasibility of integrating LLMs for multi-modal reasoning with dramatically improved performance, highlighting the great potential for LLM-assisted LVLM reasoning in future research.
6 Limitations
In this section, we analyze the limitations of the proposed method based on typical errors made by GPT-4o-mini-driven ProReason, as exemplified in Figures 6 and 7, to gain further understanding and identify potential research directions.
Cumulative Errors.
As illustrated in Figure 6, the vision expert mistakenly perceives the clock as 6:25, which misguides the reasoning of subsequent agents and ultimately leads to an incorrect conclusion. More broadly, similar misperceptions occur frequently in errors made by ProReason. This indicates that, with the assistance of LLMs, ProReason has effectively addressed the reasoning deficiencies in multi-modal tasks, while the vision expert plays a significant role for further improvement of multi-modal capabilities.
Contradictory Information among Agents.
Considering that multiple agents are engaged in the answering process, we try to find instances where contradictory information is provided by different agents, especially the vision expert and reasoning expert. However, as shown in Figure 6 and 7, when one agent (e.g., the vision expert) makes an error and the referee even hints at a possible mistake, other agents (e.g., the reasoning expert) tend to adhere to the available information instead of questioning it. This tendency results in a failure to find cases with contradictory information, and also highlights the importance of a reflection mechanism (Ji et al., 2023) in agent collaboration, which is left for future exploration.
7 Ethics Statement
We adhere strictly to the ACL Code of Ethics throughout our research. To our knowledge, the methods we introduce pose no foreseeable risks. We provide comprehensive details of the computing infrastructure used for all computational experiments in the paper, along with transparent statistics on our results and a detailed configuration of our experimental setup, including the optimal hyperparameter values. Furthermore, we will release the code upon publication to facilitate easy public reproducibility.
Appendix A Appendix
A.1 Related work
Large Visual-Language Model.
Recently, large vision-language models (LVLMs) (Bai et al., 2023; Chen et al., 2023; Liu et al., 2024b) have garnered widespread attention and demonstrated remarkable advancements in understanding and generating multi-modal contents. In the open-source domain, numerous LVLMs, like LLaVA (Liu et al., 2023c, b, 2024a; Li et al., 2024a, b) and InternVL (Chen et al., 2024b) families, have been extensively developed. In the closed-source domain, proprietary models such as GPT-4 (Achiam et al., 2023) and Gemini Pro 1.5 (Reid et al., 2024) have also achieved significantly success. Despite these advancements, existing LVLMs still encounter challenges in effectively integrating visual understanding with textual reasoning capabilities simultaneously. This limitation is particularly evident in their diminished attention to image content during visual reasoning process, such as chart interpretation and visual math reasoning, leading to degraded performance (Ghosh et al., 2024) and motivating more effective solutions.
Passive Visual Reasoning.
Extracting information from images into text can effectively assist LVLM in performing visual reasoning tasks. Visual Description Grounded Decoding (VDGD) (Ghosh et al., 2024) first describes the image before prefixing this description to the prompt, assisting LVLMs on visual reasoning tasks. Furthermore, Compositional Chain-of-Thought (CCoT) (Mitra et al., 2024) directs LVLMs to generate scene graphs (SGs) that serve as a bridge between the visual and textual domains, aiding LVLMs in subsequent tasks. However, most of these methods employ a question-agnostic and reasoning-free visual extraction process, where image descriptions are not tailored to specific questions, and no reasoning is applied to infer additional information for improved descriptions. These “passive” approaches lead to the inclusion of irrelevant or redundant information, ultimately degrading performance. In contrast, ProReason adopts question-oriented agents to collect necessary and sufficient information, effectively circumventing these drawbacks.
Multi-step Reasoning Framework.
Multi-step reasoning frameworks have been developed for LLMs to achieve better performance by breaking down complex questions into easier ones (Pan et al., 2024). As a representative method, Chain-of-Thought (CoT) (Wei et al., 2022) enhances the arithmetic and commonsense reasoning capabilities by explicitly generating intermediate reasoning steps before concluding the final answers. Tree-of-Thoughts (ToT) (Yao et al., 2024) further refines the CoT mechanism by allowing LLMs to consider multiple reasoning paths and do self-assessment before making decisions. Considering that the inherent knowledge of LLMs may not be sufficient to complete tasks, ReAct (Yao et al., 2022) integrates information retrieval into the reasoning chains, enabling models to pause to verify results and determine whether additional information is needed before proceeding. Nevertheless, multi-step reasoning frameworks designed for text-only LLMs are not fully applicable to visual reasoning tasks, and may even impair the performance of LVLMs (Ghosh et al., 2024).
A.2 Dataset
To thoroughly assess the performance of our framework, we have carried out experimental evaluations using four benchmark datasets: the Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023a), the cross-disciplinary Massive Multi-modal Understanding and Reasoning benchmark (MMMU) (Yue et al., 2023), the visual mathematical reasoning assessment MathVista (Wang et al., 2024), and the multimodal illusion detection benchmark HallusionBench (Liu et al., 2023a). Each of these benchmarks necessitates strong visual reasoning capabilities for successful task completion, and we provide concise descriptions below:
  • MME is an inclusive benchmark that encompasses 14 subtasks, designed to evaluate perceptual and cognitive abilities. Given our focus on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning.
  • MMMU evaluates multi-modal models with multidisciplinary tasks that require college-level domain-specific knowledge and detailed reasoning. It comprises 11,500 questions across 30 disciplines and 183 sub-fields, emphasizing advanced perception and domain-specific reasoning.
  • MathVista focuses on more challenging mathematical reasoning tasks that demand precise visual recognition and compositional reasoning. It includes 6,141 examples from 31 multi-modal mathematics datasets.
  • HallusionBench evaluates models’ ability to reason with images such as statistical charts, emphasizing nuanced visual understanding. It consists of 346 images paired with 1,129 questions, meticulously crafted by experts.
A.3 Supplementary Results and Analysis
A.3.1 Which One is More Crucial: Visual Understanding or Textual Reasoning?
ProReason effectively decouples the visual understanding and textual reasoning capabilities of LVLMs. However, it remains unclear which of these two capacities is more critical for visual reasoning tasks. To answer this question, we conduct comparative experiments of the following three scenarios:
  1. Llama3-LLaVA-NeXT-8B as All Sub-Agents. All sub-agents within ProReason framework are performed by Llama3-LLaVA-NeXT-8B model.
  2. GPT-4o-mini as Vision Expert. Based on the above scenario, we implement the Vision Expert with GPT-4o-mini, while keep the other textual sub-agents unchanged.
  3. GPT-4o-mini as Textual Sub-Agents. Reversely, we utilize Llama3-LLaVA-NeXT- 8B as the Vision Expert, and GPT-4o-mini for the other vision-irrelevant sub-agents.
Textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks, although both are important. As shown in Table 3, replacing either the Vision Expert or the other agents with the more capable GPT-4o-mini achieves consistent performance enhancement, highlighting the significance of both capabilities. However, substituting the textual sub-agents with GPT-4o-mini results in a more substantial performance boost compared to replacing the Vision Expert. This underscores the greater importance of textual reasoning over visual understanding for multimodal reasoning tasks, aligning with our previous analysis in Sec. 4.3 that identifies the Summarizer and Referee as the most crucial sub-agents.
Table 3: Performance of ProReason with different configurations for the relative importance assessment between visual understanding and textual reasoning capabilities on visual reasoning tasks. The red text highlights the performance improvements brought about by the introduction of GPT-4o-mini.
Model Agent Dataset
MME MMMU
GPT-4o -mini Textual Sub-Agents 84.7 (+13.2) 54.5 (+2.0)
Vision Expert 77.8 (+6.3) 53.4 (+0.9)
Llama3- LLaVA- NeXT-8B All Sub-Agents 71.5 52.5
COT 58.8 41.5
Direct 61.5 41.8
A.3.2 Referee’s Dispel of Hallucinations
Following the implementation details outlined in Section 4.1, we evaluate the ProReason system powered by GPT-4o-mini on both MMMU and HallusionBench under varying attempt allowances (i.e., 1, 3, and 5 attempts). Every unsuccessful attempt reflects the Referee’s persistent determination that the information stored in Memory is insufficient to solve the problem. Before each new attempt, Memory is systematically cleared, ensuring the removal of information deemed irrelevant by the Referee’s assessment. Consequently, increased attempt allowances essentially empower the Referee with enhanced opportunities for information filtration. The observed performance variations of ProReason across different attempt quotas demonstrate the critical impact of Referee’s decision-making mechanism and information filtering efficacy on system capability.
The Referee module effectively filters hallucinated information to enhance the visual comprehension capabilities of our framework. As demonstrated in Table 4, the performance improvement on HallusionBench (2.8%) significantly outpaces that on MMMU (1.5%) as attempt opportunities increase from 1 to 5. Given HallusionBench’s dual emphasis on reasoning proficiency and precise evaluation of visual details/hallucination control compared to MMMU, these results suggest that the Referee mechanism can effectively identify erroneous or irrelevant visual information, thereby strengthening ProReason’s capacity for meticulous visual understanding.
Table 4: Impact on HallusionBench and MMMU performance across different attempt allowances.
Dataset Attempt Allowances
1 3 5
MMMU 60.1 59.9 61.6
HallusionBench 57.1 58.9 59.9
A.3.3 Performance of ProReason on VQA task
ProReason also excels in conventional VQA tasks. We perform an evaluation on the VQA benchmark A-OKVQA dataset (Schwenk et al., 2022). The visual complexity of A-OKVQA is comparable to MME, but its reasoning requirements are less demanding. Table 5 illustrates the performance of ProReason, powered by GPT-4o-mini, on the A-OKVQA, MME, and MMMU benchmarks. ProReason achieves the highest performance across all three benchmarks, indicating the positive impact of ProReason on any form of reasoning-involved multi-modal tasks.
ProReason is particularly effective on datasets with greater complexity due to its decoupled capabilities and LLM-assisted reasoning. Table 5 demonstrates that ProReason outperforms the "Direct" method by 2.7%, 11.2%, and 13.2% on A-OKVQA, MME, and MMMU, respectively, as the visual and reasoning complexity increases. This underscores the advantage of ProReason in handling more challenging tasks, making it a robust choice for complex problem-solving scenarios.
Table 5: Performance of multiple approaches with GPT-4o-mini across three visual benchmarks.
GPT-4o-mini Dataset
A-OKVQA MME MMMU
Direct 78.9 79.2 48.4
VDGD 79.4 82.3 51.4
CCoT 79.2 80.8 54.2
COT 80.9 87.8 58.5
ReAct 80.6 87.3 54.8
ProReason 81.3 (+2.7) 90.4 (+11.2) 61.6 (+13.2)
A.3.4 Evaluation of ProReason based on LLMs
In Sec. 2.1, we utilize a pipeline based on LLMs to evaluate the correlation between model responses and image content. We aim to apply this effective method once more to assess the performance of ProReason. Initially, we adopted the reasoning chains generated by GPT-4o-mini (CoT) as the standard answers, given its superior performance with a score of 58.5 on the MMMU benchmark, compared to other models. However, due to ProReason’s enhancement of GPT-4o-mini’s score to 61.6, the use of GPT-4o-mini’s answers became inadequate. Consequently, we opted for GPT-4o to generate the standard answers, achieving a score of 69.1 on MMMU.
As shown in Table 6, we evaluate the reasoning process of GPT-4o-mini (CoT), the Memory of ProReason (ProReason-Memory), and the reasoning process of Summarizer (ProReason-Summarizer), which generates the final answer of ProReason with the prompt template in Figure 12. This assessment focuses on three key metrics: the relevance to standard answers (RE ↑), the degree of redundant information (RI ↓), and the extent of missing information (MI↓), where arrows indicate the directions of improvement.
Specifically, compared to GPT-4o-mini (CoT), ProReason-Summarizer produces more relevant answers with less redundancy and deficiency, aligning with its improved performance. Compared to ProReason-Summarizer, ProReason-Memory exhibits the same RE, higher RI and lower MI scores. This suggests that ProReason allows some redundancy to prevent information loss in memory, as the former typically leads to more serious consequences than the latter. Subsequently, Summarizer can leverage its powerful reasoning capabilities to select the most relevant memory.
Table 6: Performance of ProReason driven by GPT-4o-mini assessed by LLMs compared to CoT on MMMU benchmark. Mainly includes three key metrics: the relevance to standard answers (RE ↑), the degree of redundant information (RI ↓), and the extent of missing information (MI ↓), where arrows indicate the directions of improvement.
GPT-4o-mini Metrics
RE ↑ RI ↓ MI ↓
COT 4.67 3.33 1.40
ProReason-Memory 4.83 3.66 1.17
ProReason-Summarizer 4.83 2.88 1.33
A.3.5 Frequency of selection of various experts
As listed in Table 7, we evaluate the frequency of the Dispatcher choosing the Vision Expert or Reasoning Expert on both MME and MMMU benchmarks, with MMMU requiring higher visual and reasoning abilities. Specifically, compared to MME, the frequencies for both the Vision and Reasoning Experts are higher on the MMMU benchmark, aligning with their difficulty levels. Together with the results in Table 1 of our submission, ProReason can adaptatively increase the frequencies of experts, and provide consistent performance improvements (i.e., 11.2% and 13.2%). Despite the lower frequency of the reasoning expert, the significant performance enhancement highlights the importance of LLM-assisted reasoning capabilities for reasoning-essential questions. Additionally, the frequency of the Vision expert exceeding 1 underscores the importance of referees, which controls the loop to call experts multiple times, alleviating the issue of insufficient information.
Table 7: Frequency of the Dispatcher choosing the Vision Expert or Reasoning Expert on both MME and MMMU benchmarks.
Dataset GPT-4o-mini
Vision Expert Reasoning Expert
MME 1.16 0.12
MMMU 1.64 0.38
Table 8: Performance of three recent LVLMs on MMMU dataset with different assisting techniques.
Model Method
Direct CoT VDGD
Llama3-LLaVA-NeXT-8B 41.8 41.5 42.3
LLaVA-OneVision-Qwen2-7B-OV 49.1 46.5 49.7
Qwen-VL-Chat 36.2 35.4 37.1
Table 9: Image Relevance Score of Chain-of-Thought reasoning traces for “True” and “False” responses of three LVLMs, respectively.
Model Image Relevance Score
True False
Llama3-LLaVA-NeXT-8B 3.70 3.36
LLaVA-OneVision-Qwen2-7B-OV 4.27 3.67
Qwen-VL-Chat 3.63 2.80
Table 10: Effectiveness evaluation of passive captions along Detail Level, Question Relevance, and Reasoning Effective Info Inclusion. “True” and “False” denote the response correctness of Llama3-LLaVA-NeXT-8B.
Score Llama3-LLaVA-NeXT-8B
True False
Detail Level 4.43 3.93
Question Relevance 3.87 3.30
Reasoning Effective Info Inclusion 3.91 3.57
Table 11: Average token consumption of ProReason and ReAct with GPT-4o-mini model on the MME and MathVista benchmarks.
Method MME MathVista
Input Output Input Output
ProReason 1286.8 327.2 2238.6 788.6
ReACT 1645.0 197.0 3092.8 845.1
Table 12: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks. For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini. The blue text indicates the performance decline compared to the scenario with all agents performed by GPT-4o-mini.
Model Agent Dataset
MME MMMU
GPT-4o-mini 90.4 61.6
Llama3- LLaVA- NeXT-8B Dispatcher 88.8 (-1.6) 60.9 (-0.7)
Vision Expert 84.7 (-5.7) 54.5 (-7.1)
Reasoning Expert 88.7 (-1.7) 60.2 (-1.4)
Referee 89.6 (-1.1) 51.5 (-10.1)
Summarizer 84.2 (-6.2) 51.0 (-10.6)
Table 13: Impact of decoupling visual perception and textual reasoning on ProReason performance with results shown from sub-agent consolidation experiments. The blue text highlights the performance degradation due to sub-agent integration.
GPT-4o-mini Dataset
MMMU MME
ProReason 61.6 91.9
CoT 58.5 87.8
Vision & Reasoning
Expert Integration
58.6 (-3.0) 90.4 (-1.5)
Dispatcher,
Vision & Reasoning Expert,
& Referee Integration
57.8 (-3.8) 90.8 (-1.1)
All Sub-Agents Integration 56.8 (-4.8) 89.2 (-2.7)
Table 14: Performance of multiple approaches with two base models across four visual reasoning benchmarks. Different from Table 1, MME scores are calculated with the official method here.
Model Method Dataset
MME MMMU MathVista HallusionBench
Llama3-LLaVA- NeXT-8B Direct 35.9 41.8 37.1 45.8
VDGD 37.4 42.3 36.1 44.2
CCoT 35.4 40.5 36.8 37.4
CoT 25.8 41.5 35.9 43.1
ReAct 38.5 46.7 31.7 43.6
ProReason 41.6 50.5 38.8 50.9
GPT-4o-mini Direct 55.2 48.4 53.0 56.0
VDGD 58.4 51.4 51.2 52.4
CCoT 53.5 54.2 53.6 56.7
CoT 72.9 58.5 53.8 56.3
ReAct 71.2 54.8 49.3 51.1
ProReason 77.9 61.6 54.9 59.9
Table 15: Performance of Llama3-LLaVA-NeXT-8B Vision Expert, assisted by different text-only LLMs, on various benchmarks under ReAct and ProReason frameworks. Different from Table 2, MME scores are calculated with the official method here.
Vision Expert Textual Sub-Agents Method Dataset
MME MMMU MathVista Hallu.
Llama3- LLaVA- NeXT-8B Llama3-LLaVA- NeXT-8B ReAct 385 46.7 31.7 43.6
ProReason 416 50.5 38.8 50.9
Qwen2.5-72B- Instruct ReAct 428 50.4 34.6 40.4
ProReason 603 56.8 48.8 52.3
GPT-4o-mini ReAct 473 48.4 36.2 46.7
ProReason 613 54.5 41.7 53.1
Table 16: Performance of ProReason with different configurations for the relative importance assessment between visual understanding and textual reasoning capabilities on visual reasoning tasks. Different from Table 3, MME scores are calculated with the official method here.
Dataset GPT-4o-mini Llama3-LLaVA-NeXT-8B
Textual Sub-Agents Vision Expert All Sub-Agents COT Direct
MME 613 512 416 286 359
MMMU 54.5 53.4 52.5 41.5 41.8
Table 17: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks. For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini. Different from Table 13, MME scores are calculated with the official method here.
Model Agent Dataset
MME MMMU
GPT-4o-mini 779 61.6
Llama3- LLaVA- NeXT-8B Dispatcher 746 60.9
Vision Expert 613 54.5
Reasoning Expert 743 60.2
Referee 762 51.5
Summarizer 641 51.0
Table 18: Impact of decoupling visual perception and textual reasoning on ProReason performance with results shown from sub-agent consolidation experiments. Different from Table 13, MME scores are calculated with the official method here.
GPT-4o-mini Dataset
MMMU MME
ProReason 61.6 779
CoT 58.5 729
Vision & Reasoning
Expert Integration
58.6 673
Dispatcher,
Vision & Reasoning Expert,
& Referee Integration
57.8 714
All Sub-Agents Integration 56.8 711
A.4 Demonstrative Examples
Figure 2: An example with three reasoning frameworks: fine-grained caption, chain-of-thought, and ProReason. ProReason enables LVLMs to proactively acquire necessary information in a question-oriented manner, and predicts answers based on the collected information. Apparently, ProReason is superior to previous methods, which often describe question-irrelevant visual details, or overlook informative elements. Green indicates correct information or conclusions, while red signifies incorrect ones.
Figure 3: A complete reasoning process of ProReason for the case shown in Figure 2.
Figure 4: Additional examples of images-unrelated Chain-of-Thought reasoning.
Figure 5: Additional examples of question-unrelated passive image captions.
Figure 6: A typical mistake made by ProReason. The vision expert incorrectly identifies 4:30 as 6:25, leading the other agents to base their judgments on this erroneous information, and ultimately resulting in the wrong conclusion.
Figure 7: A typical mistake made by ProReason. The vision expert incorrectly provided the coordinates of the three vertices of the triangle, leading the other agents to base their judgments on this erroneous information, and ultimately resulting in the wrong conclusion.
A.5 Prompts
Figure 8: Prompt templates of Relevance Score and caption effectiveness evaluation.
Figure 9: Prompt templates of Chain-of-Thought, fine-grained image captions generation, Compositional Chain-of-Thought (CCoT), and ReAct.
Figure 10: Prompt templates of Dispatcher.
Figure 11: Prompt templates of Vision Expert, Reasoning Expert, Referee, and Summarizer.
Figure 12: Prompt templates of LLMs-based ProReason evaluation.
Figure 13: Prompt templates of Vision and Reasoning Expert Integration.
Figure 14: Prompt templates of Dispatcher, Vision Expert Reasoning Expert and Referee Integration.
Figure 15: Prompt templates of All Five Sub-Agent Integration.