### v3

# ProReason: Multi-Modal Proactive Reasoning with Decoupled Eyesight and Wisdom

Abstract
Large vision-language models (LVLMs) have witnessed significant progress on visual understanding tasks. However, they often prioritize language knowledge over image information on visual reasoning tasks, incurring performance degradation. To tackle this issue, we first identify the drawbacks of existing solutions (i.e., limited multi-modal reasoning capacities, and insufficient and irrelevant visual descriptions). We then decompose visual reasoning process into two stages: proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), and introduce a novel visual reasoning framework named ProReason. This framework features decoupled vision-reasoning capabilities and multi-run proactive perception. Briefly, given a multi-modal question, ProReason iterates proactive information collection and reasoning until the answer can be concluded with necessary and sufficient visual descriptions. Notably, the disassociation of capabilities allows seamless integration of existing large language models (LLMs) to compensate for the reasoning deficits of LVLMs. Our extensive experiments demonstrate that ProReason outperforms existing multi-step reasoning frameworks on various benchmarks for both open-source and closed-source models, with the average performance gain reaching 13.2%. Besides, the integration of LLMs allows ProReason to produce high-quality visual reasoning data, which empowers ProReason-distilled models (i.e., ProReason-VL and ProReason-Q3) to achieve superior performance in downstream tasks. Our insights into existing solutions and the decoupled perspective for feasible integration of LLMs illuminate future research on visual reasoning techniques, especially LLM-assisted ones.
ProReason: Multi-Modal Proactive Reasoning
with Decoupled Eyesight and Wisdom
Jingqi Zhou, Sheng Wang, Jingwei Dong, Kai Lui, Lei Li, Jiahui Gao, Jiyue Jiang, Lingpeng Kong, Chuan Wu The University of Hong Kong, The Chinese University of Hong Kong, Xi’an Jiaotong University {u3011211, u3009618}@connect.hku.hk, dongjingwei@stu.xjtu.edu.cn, u3638070@connect.hku.hk, nlp.lilei@gmail.com, sumiler@connect.hku.hk, jiangjy@link.cuhk.edu.hk, {lpk, cwu}@cs.hku.hk
1 Introduction
In recent years, large language models (LLMs) (Yang et al., 2024; Dubey et al., 2024; Team et al., 2023; Jiang et al., 2023) have experienced explosive growth in their capabilities, driving significant advancements across various fields (Shao et al., 2023; Guo et al., 2024; Shao et al., 2024). This progress has also sparked interest in developing large vision-language models (LVLMs) (Bai et al., 2025; Chen et al., 2024b; Bai et al., 2023), which, like LLaVA (Li et al., 2024b), have achieved remarkable performance in multi-modal understanding tasks. However, state-of-the-art (SOTA) LVLMs still struggle to handle visual understanding with textual reasoning simultaneously due to inherent modality differences. For example, Ghosh et al. (2024) demonstrate that LVLMs often rely more on their prior language knowledge, neglecting visual information in multi-modal reasoning tasks, such as visual chart understanding and math reasoning, resulting in performance degradation. Figure 2.b illustrates a typical case of this issue, where the reasoning process remains irrelevant to the image.
To address the challenges, several visual reasoning frameworks have been proposed. Specifically, Ghosh et al. (2024) and Mitra et al. (2024) convert visual information in images into textual descriptions to aid LVLMs in reasoning. However, their visual extraction process is not targeted at a given question (i.e., question-agnostic), termed as "passive", and omits reasoning mechanisms to infer extra information for better descriptions(i.e., reasoning-free). These limitations result in irrelevant or inadequate information, ultimately degrading performance. Furthermore, these frameworks are powered by a single LVLM, leading to a reasoning process that conflates visual understanding with textual reasoning abilities, failing to mitigate the challenge faced by LVLMs in effectively managing both capabilities.
To resolve these problems, we propose ProReason, a multi-modal reasoning framework featuring decoupled vision-reasoning capabilities. As illustrated in Figure 1, we decouple multi-modal reasoning capacity into two sub-tasks: proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom). The former extracts visual information in a question-oriented and reasoning-involved manner, while the latter integrates all information to draw final conclusions. Specifically, during the visual perception stage, a Dispatcher first selectively engages a Vision Expert to capture additional visual information, or an Insight Expert to derive intermediate inferences. A Referee then determines whether sufficient information is gathered to proceed to the reasoning stage, where a Summarizer produces the final answer. Unlike passive methods, all sub-agents operate based on the given question and known information, effectively avoiding irrelevant information redundancy or insufficiency. Notably, decoupled vision-reasoning eliminates the need for LVLMs to handle vision-irrelevant roles (i.e., Dispatcher, Insight Expert, Referee, and Summarizer), enabling seamless integration of existing LLMs with proven strong reasoning abilities (Chang et al., 2024), thereby alleviate the limitations of LVLMs. In addition, the high-quality reasoning data generated by LLM-assisted ProReason can be effectively distilled into downstream models for inherent performance improvement.
Empirically, we evaluate ProReason across multiple challenging visual reasoning benchmarks with both open-source and closed-source models. Extensive experiments demonstrate that ProReason exhibits significant advantages in two key aspects: (1) As a visual reasoning framework, ProReason achieves consistent and substantial performance improvements across multiple benchmarks, with the average performance gain reaching 13.2%, validating the effectiveness of its decoupled vision-reasoning architecture and proactive visual feature extraction mechanism; (2) ProReason effectively integrates existing LLMs to generate high-quality visual reasoning process, empowering ProReason-distilled models (i.e., ProReason-VL and ProReason-Q3) with superior visual reasoning capabilities. The above results, coupled with the ablation study in Section 4.5, demonstrate the substantial advantages of decoupled vision-reasoning, while highlighting the potential of LLM-assisted LVLM reasoning and distillation strategies.
The main contributions of this work are threefold:
  • We propose a novel multi-modal reasoning framework named ProReason, featuring decoupled vision-reasoning and iterative proactive perception capabilities, effectively mitigating the drawbacks of previous methods.
  • Extensive experiments consistently highlight the significant superiority of ProReason and necessity of each component across multiple visual reasoning tasks and model series, illuminating the great potential of LLM-assisted LVLM reasoning.
  • ProReason-distilled models also exhibits remarkable enhancements over vanilla counterparts, showcasing the feasibility of LLM-assisted LVLM improvement in the future.
2 Preliminary Observations
Ghosh et al. (2024) demonstrate that the limited multimodal reasoning abilities of LVLMs lead to an overreliance on linguistic priors, thus neglecting visual inputs and ultimately degrading their performance. Their Visual Description Grounded Decoding (VDGD) mitigates visual oversight by converting images into comprehensive textual descriptions to inform reasoning processes. However, such passive visual reasoning techniques suffer insufficient and irrelevant visual information. To support this claim, we generate fine-grained image captions using GPT-4o-mini [^1] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence with the prompt shown in Figure 9. We then incorporate these captions into the prompts for LVLMs to facilitate the reasoning process. We analyze the performance of this approach on the challenging multi-modal MMMU dataset, which requires college-level knowledge and fine-grained reasoning, using recent open-source LVLMs listed in Section 4.1 As shown in Table 4, while these image descriptions improve the performance of LVLMs, the gains are marginal, consistently amounting to less than 1%. This underscores the limited utility of captions generated by passive methods.
For further demonstration, inspired by Liu et al. (2023c), we use the prompt instructions in Figure 8 to instruct GPT-4 to analyze the generated captions along three dimensions: Detail Level, Question Relevance, and Reasoning Effective Info Inclusion, measuring the richness of detail, relevance to the given question, and the inclusion of information that is essential for reasoning, respectively. Meanwhile, since the reasoning process of GPT-4o-mini on MMMU contains key information necessary for solving the problems, we use it as a reference answer to aid evaluation. As shown in Table 5, the captions for correct responses of Llama3-LLaVA-NeXT-8B receive higher scores across all three criteria, highlighting the importance of better captions for multi-modal reasoning. Additionally, all captions score significantly lower in the Question Relevance and Reasoning Effective Info Inclusion dimensions than the Detail Level dimension, indicating that while the captions are detailed, they often lack relevance to the questions. Figure 2 shows a case where Llama3-LLaVA-NeXT-8B utilizes fine-grained captions to solve a question from the MMMU benchmark. As illustrated, although the caption exhaustively describes the image content, it incorrectly describes the wires in the image as octagons, and misses information about the locations of these wires. This information is irrelevant to the target question, thus offering minimal assistance to LVLMs. In summary, our analysis highlights the drawbacks of passive visual reasoning enhancement techniques in terms of information insufficiency and redundancy, due to their question-agnostic property.
Figure 1: Overview and comparison of ProReason, VDGD and ReAct. Unlike existing works (e.g., VDGD and ReAct), our proposed method decouples visual perception and textual reasoning while allowing the model to actively acquire necessary information from the images, achieving superior performance.
(1) ProReason
(2) Passive Visual Reasoning Enhancement Technique
(3) Multi-step Reasoning Framework for LLMs
3 Method
As depicted in Figure 1 and exemplified in Figure 3, ProReason presents an innovative decoupling of the visual reasoning process into two distinct phases framed through the lens of LVLM capabilities: Proactive Visual Perception (i.e., eyesight) and Textual Reasoning (i.e., wisdom). The entire workflow consists of five functionally distinct yet inter-cooperative sub-agents, along with a Memory component, facilitating multi-modal reasoning performance.
3.1 Proactive Visual Perception
Proactive Visual Perception is the core of question-oriented visual information extraction, driven by four sub-agents: Dispatcher, Vision Expert, Insight Expert and Referee. The Dispatcher breaks down the original question, selectively directing the Vision Expert to capture specific visual information, or instructing the Insight Expert to analyze known information to derive more. The responses from both experts are stored in a textual Memory component. The Referee then evaluates whether the information stored in Memory is sufficient to answer the original question.
Formally, given an image $I$ and its corresponding textual question $Q$, the Dispatcher decides to consult the Vision Expert or Insight Expert, based on the analysis of $Q$ and the known information in the Memory (if not empty). The Dispatcher then generates a query $q$ for the chosen expert. If the Vision Expert is selected, it takes the image $I$ and query $q$ as input, and generates an answer $A_{vision}$, which is then stored in the Memory. When the Insight Expert is selected, it provides a response consisting of the inference process and the final answer $A_{insight}$ based on the query $q$ and known information in the Memory, before only $A_{insight}$ is stored in Memory. The Referee then evaluates the available information in the Memory concerning the question $Q$. If the Memory contains adequate information to answer the question $Q$, the Referee outputs the identifier “SOLVABLE”; otherwise, it outputs “UNSOLVABLE”. If the Referee’s output is SOLVABLE, the workflow precedes to the Textual Reasoning phase. Conversely, if the output is UNSOLVABLE, the above process will be re-executed to gather more necessary information.
In the Proactive Visual Perception phase, the Dispatcher collaborates with the Vision Expert and Insight Expert to achieve question-oriented and reasoning-involved visual information extraction, while the Referee ensures informational completeness and prevents omissions. These four sub-agents work closely together, thereby overcoming the limitations of passive methods. Notably, the Memory component allows ProReason to keep compact information, and avoids lengthy reasoning traces like CoT (Wei et al., 2022) and ReAct (Yao et al., 2022), thereby suffering less from redundant information (for a detailed analysis, refer to Section 4.4 and Table 3).
3.2 Textual Reasoning Step
The Textual Reasoning step focuses on integrating the available information in the Memory, and providing the final answer to the question $Q$. This step is mainly powered by a sub-agent called Summarizer. Once the Referee confirms that the Memory contains sufficient information to address question $Q$ and outputs the SOLVABLE identifier, the Summarizer will be invoked to perform a detailed reasoning and generate a final answer based on the question $Q$ and Memory. This final answer is then evaluated using performance metrics, while the Summarizer’s high-quality reasoning outputs can be utilized to train downstream models, enhancing their visual reasoning capabilities.
3.3 Advantages of ProReason
Decoupled Visual Reasoning.
In ProReason, the multi-modal reasoning process is decomposed into visual perception and textual reasoning stages, each executed by separate agents. These agents are then effectively organized through a designated pipeline, significantly enhancing the ability of LVLMs to tackle visual reasoning tasks.
LLM-assisted multi-modal reasoning.
Decoupled vision-reasoning eliminates the necessity for LVLMs to manage vision-irrelevant sub-agents, allowing the seamless integration of existing LLMs with established strong reasoning abilities, thereby endowing ProReason with superior visual reasoning performance.
Downstream task model enhancement.
By generating high-quality data, ProReason effectively distills its superior capabilities into downstream task models, enabling them to demonstrate exceptional performance in complex visual reasoning tasks.
Reduced information Mission or Redundancy.
Through sub-agent collaboration in Proactive Visual Perception, ProReason extracts essential visual details via question-oriented manner, preventing information omission or redundancy. Meanwhile, the Memory component retains only Vision Expert’s observations and Insight Expert’s conclusions, creating compact representation that minimizes irrelevant information. These advantages effectively reduce the overhead when ProReason functions as a reasoning framework.
4 Experiments
In this section, we first evaluate the performance of our ProReason framework against recent baselines on multiple benchmarks, followed by an in-depth ablation analysis of different components.
4.1 General Setup
Datasets.
To comprehensively validate the performance of our framework, we conduct experiments across four benchmarks: Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023)  [^2] Due to our emphasis on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning. To facilitate the comparison across different benchmarks, the results for MME benchmark are calculated by the percentage of correct answers out of the total answers. , Massive Multi-discipline Multi-modal Understanding and Reasoning (MMMU) (Yue et al., 2023), MathVista (Wang et al., 2024), and HallusionBench (Liu et al., 2023a). All of them require visual reasoning capabilities to complete the tasks correctly, and are introduced briefly in Section A.2.
Base Models.
We employ GPT-4o-mini, Llama3-LLaVA-NeXT-8B and Qwen2.5-VL-7B- Instruct (Bai et al., 2025) [^3] The model is deployed using CUDA https://developer.nvidia.com/cuda-toolkit on a NVIDIA A100 graphics cards. to drive ProReason, and Qwen2.5-72B-Instruct (Yang et al., 2024) and Qwen3-32B (Team, 2025) are selected to drive the text-only sub-agents, enabling LLM-assisted visual reasoning. In addition Qwen2.5-VL-7B-Instruct is also utilized as a downstream task model, which is trained with the data generated by ProReason. These models are chosen for their leading performance and popularity.
Baselines.
In addition to the basic method where models are instructed to answer questions directly, we compare ProReason with the following peer methods. First, to explore the effectiveness of directly migrating LLM solutions to LVLMs, we select two multi-step reasoning frameworks from LLMs: Chain of Thought (CoT) (Wei et al., 2022) and ReAct (Yao et al., 2022). Second, to evaluate the advantages of proactive information extraction in ProReason, we examine two passive visual reasoning frameworks, VDGD (Ghosh et al., 2024) and CCoT (Mitra et al., 2024), which assist LVLM reasoning by extracting image information into text. Additionally, we consider R1-Onevision-7B (Yang et al., 2025), an LVLM based on Qwen2.5-VL-7B-Instruct with deep thinking capabilities similar to Deekseek-R1 (Guo et al., 2025) and OpenAI o1444https://openai.com/index/introducing-openai-o1-preview/. We compare it with ProReason-Q3, which also exhibits deep thinking abilities. Further details on the baselines and implementation are provided in Section A.3.
Implementation Details.
The prompt templates for all methods are shown in Figures 9, 10 and 12. Specifically for ProReason, to prevent infinite loops, if the Dispatcher selects the Vision Expert or Insight Expert to obtain information up to 5 consecutive times, and the Referee still determines that the existing knowledge in Memory remains insufficient to resolve the question, the Memory will be cleared to restart the information acquisition process. If the system fail to break the loop after 5 attempts, the proactive visual perception phase will be immediately terminated. Subsequently, the Summarizer will generate the final answer based on the available information in Memory from the last attempted iteration. This setup, refined through multiple trials, is the most effective.
Table 1: Performance of multiple approaches with three base models across four visual reasoning benchmarks. “Hallu.” is the abbreviation of HallusionBench. The abbreviations 4o-mini, Qwen72B, and Qwen3 refer to GPT-4o-mini, Qwen2.5-72B-Instruct, and Qwen3-32B, respectively. "Assisted" stands for LLM-assisted reasoning, which involves replacing the textual sub-agents within frameworks with corresponding LLMs, as detailed in Table 14. Based on the performance of the direct method, red and blue signify the improvement and degradation, respectively.
Model Method Dataset
MME MMMU MathVista Hallu. Average
GPT-4o-mini Direct 79.2 48.4 53.0 56.0 59.2
VDGD 82.3 (+3.1) 51.4 (+3.0) 51.2 (-1.8) 52.4 (-3.6) 59.3 (+0.1)
CCoT 80.8 (+1.6) 54.2 (+5.8) 53.6 (+0.6) 56.7 (+0.7) 61.3 (+2.1)
CoT 87.8 (+8.6) 58.5 (+10.1) 53.8 (+0.8) 56.3 (+0.3) 64.1 (+4.9)
ReAct 87.3 (+8.1) 54.8 (+6.4) 49.3 (-3.7) 51.1 (-4.9) 60.6 (+1.4)
ProReason 91.9 (+12.7) 61.6 (+13.2) 54.9 (+1.9) 59.9 (+3.9) 67.1 (+7.9)
Llama3-LLaVA-NeXT-8B Direct 61.5 41.8 37.1 45.8 46.6
VDGD 68.8 (+7.3) 42.3 (+0.5) 36.1 (-1.0) 44.2 (-1.6) 47.8 (+1.2)
CCoT 68.9 (+7.4) 40.5 (-1.3) 36.8 (-0.3) 37.4 (-8.4) 45.9 (-0.7)
CoT 58.8 (-2.7) 41.5 (-0.3) 35.9 (-1.2) 43.1 (-2.7) 44.8 (-1.8)
ReAct 68.5 (+7.0) 46.7 (+4.9) 31.7 (-5.4) 43.6 (-2.2) 47.6 (+1.0)
+ 4o-mini Assisted 73.6 (+12.1) 48.4 (+6.6) 36.2 (-0.9) 46.7 (+0.9) 51.2 (+4.6)
+ Qwen72B Assisted 71.0 (+9.5) 50.4 (+8.6) 34.6 (-2.5) 40.4 (-5.4) 49.1 (+2.5)
ProReason 71.5 (+10.0) 50.5 (+8.7) 38.8 (+1.7) 50.9 (+5.1) 52.9 (+6.3)
+ 4o-mini Assisted 84.7 (+23.2) 54.5 (+12.7) 41.7 (+4.6) 53.1 (+7.3) 58.5 (+11.9)
+ Qwen72B Assisted 81.3 (+19.8) 56.8 (+15.0) 48.8 (+11.7) 52.3 (+6.5) 59.8 (+13.2)
Qwen2.5-VL-7B-Instruct Direct 74.2 51.8 63.3 53.8 60.8
VDGD 74.6 (+0.4) 52.3 (+0.5) 62.1 (-1.2) 53.9 (+0.1) 60.7 (-0.1)
CCoT 82.7 (+8.5) 52.2 (+0.4) 61.7 (-1.6) 55.8 (+2.0) 63.1 (+2.3)
CoT 72.3 (-1.9) 53.6 (+1.8) 63.7 (+0.4) 55.9 (+2.1) 61.4 (+0.6)
ReAct 81.9 (+7.7) 51.6 (-0.2) 60.4 (-2.9) 52.9 (-0.9) 61.7 (+0.9)
+ Qwen72B Assisted 83.1 (+8.9) 52.5 (+0.7) 62.2 (-1.1) 54.5 (+0.7) 63.1 (+2.3)
ProReason 83.8 (+9.6) 57.0 (+5.2) 63.2 (-0.1) 56.4 (+2.6) 64.6 (+3.8)
+ Qwen72B Assisted 92.7 (+18.5) 64.6 (+12.8) 64.0 (+0.7) 60.6 (+6.8) 70.5 (+9.7)
+ Qwen3 Assisted 90.7 (+16.5) 66.2 (+14.4) 67.2 (+3.9) 59.8 (+6.0) 71.0 (+10.2)
4.2 Main Results
ProReason exhibits significant and consistent performance enhancement over baselines across all the benchmarks. As listed in Table 1, despite better performance than the direct method on MME dataset, VDGD and CCoT fail to demonstrate consistent improvements on the other datasets. In contrast, ProReason consistently surpasses all other baselines across all benchmarks for every base model, enhancing the average performance of GPT-4o-mini by 7.9%, demonstrating the superiority and task robustness of ProReason. Section A.5.7 and Table 13 evaluate ProReason on additional benchmarks, including reasoning and VQA tasks, further proving its strong performance and effectiveness across various multi-modal tasks.
Decoupling the visual perception and textual reasoning capabilities of an LVLM outperforms their simultaneous inherent usage. Table 1 illustrates that CoT, utilizing both capabilities simultaneously, does not exhibit consistent performance enhancements over the "Direct" method when applied to the Llama3-LLaVA-NeXT-8B and Qwen2.5-VL-7B-Instruct models. In contrast, despite the same models, ProReason alternates between visual information acquisition and textual reasoning processes, allowing to leverage each capability more effectively. This enables ProReason to consistently outperform CoT with both Llama3-LLaVA-NeXT-8B and GPT-4o-mini across all benchmarks, validating the effectiveness of capability decoupling.
Proactive information acquisition surpasses peer passive methods, especially in complex visual reasoning tasks. Specifically, compared to MME, MathVista and HallusionBench present higher image complexity and question difficulty, and thus require stronger visual understanding and textual reasoning capabilities. This leads to performance degradation of passive methods (i.e., VDGD and CCoT), highlighting their limited applicability to complex visual reasoning tasks. In contrast, ProReason achieves notable performance improvements, up to 5.1%, by proactively acquiring visual information from images rather than generating question-agnostic captions. This aligns with out previous obervations in Section 2 that passive methods introduce substantial information redundancy or omission, misleading subsequent reasoning processes.
Text-only LLMs can be effectively integrated into ProReason for dramatically enhanced performance. As mentioned in Section 3.3, the decoupled visual perception and textual reasoning capabilities facilitate the seamless integration of text-only LLMs. To demonstrate the utility of this advantage, we fix the Vision Expert and replace the textual sub-agents in ProReason with text-only LLMs, according to the configuration in Table 14. As listed in Table 1, with the assistance of powerful existing LLMs, the Llama3-LLaVA-NeXT-8B Vision Expert receives remarkable performance boost across all benchmarks, particularly by 15% on MMMU and 11.7% on MathVista, compared to directly providing answers. In addition, by configuring the Summarizer to Qwen3-32B, ProReason acquires the same kind of deep reasoning capability as Deekseek-R1 and OpenAI o1, resulting in a 10.2% average performance improvement for the Qwen2.5-VL-7B-Instruct Vision Expert. In contrast, ReAct gains a much smaller improvement. This highlights the unique advantage of ProReason in leveraging existing text-only LLMs for enhanced performance. Notably, this advantage may open new avenues for continuously pushing the performance limits of LVLMs with the assistance of existing powerful LLMs.
Table 2: Performance of different models across two visual benchmarks. Based on the performance of the base model(i.e., Qwen2.5-VL-7B-Instruct), red and blue signify the improvement and degradation, respectively.
Dataset Model
Qwen2.5-VL-7B-Instruct R1-Onevision-7B ProReason-VL ProReason-Q3
MME 72.3 88.1 (+15.8) 91.2 (+18.9) 90.8 (+18.5)
MMMU(val.) 53.6 53.1 (-0.5) 65.4 (+11.8) 67.4 (+13.8)
Average 63.0 70.6 (+7.6) 78.3 (+15.3) 79.1 (+16.1)
Table 3: Average token and time consumption of multiple approaches with GPT-4o-mini model on the MME and MathVista benchmarks.
Dataset Method GPT-4o-mini
Direct VDGD CCoT CoT ReAct ProReason
MME Input 393.9 1020.9 1024.3 403.9 1645.0 1286.8
Output 5.9 155.6 254.3 103.4 197.0 327.2
Time(s) 6.1 12.3 12.8 7.3 18.9 18.4
MathVista Input 368.4 955.5 961.2 375.4 3092.8 2238.6
Output 51.8 263.3 307.3 479.7 845.1 788.6
Time(s) 4.2 11.5 12.3 12.0 28.8 24.8
4.3 Downstream Task Model Enhancement
High-quality visual reasoning data produced by ProReason significantly improves the performance of downstream task models. We select the test set that has no answer in the MMMU dataset and gather reasoning processes paired with corresponding answers on this set, generated by two configurations detailed in Table 14: ProReason + Qwen72B Assisted and ProReason + Qwen3 Assisted. After eliminating samples with inconsistent answers between these two configurations, we obtain two filtered datasets each containing 5,980 entries. These datasets are then used to fine-tune Qwen2.5-VL-7B-Instruct (Section A.4), yielding two models: ProReason-VL and ProReason-Q3. Notably, ProReason-Q3 inherits the deep reasoning capabilities of Qwen3-32B, which incorporates a reasoning process containing a <think>... </think> part.
Table 2 illustrates that while R1-Onevision-7B features deep reasoning capabilities and shows performance enhancements on MME, it fails to demonstrate improvement on the MMMU validation set. In contrast, both ProReason-VL and ProReason-Q3 achieve performance gains exceeding 10% on the MMMU validation set and also exhibit significant improvements on MME. This indicates that ProReason effectively transfers the exceptional reasoning capabilities of LLM-integrated decoupled systems to downstream task models, highlighting the potential for leveraging the robust abilities of existing LLMs to continuously enhance the performance of LVLMs.
4.4 Efficiency and Complexity Analysis
ProReason improves both efficiency and performance compared to the baseline method. Table 3 presents the evaluation of average token consumption and time expenditure of different methods on the MME and MathVista datasets. Notably, ProReason requires significantly fewer tokens than ReAct, while achieving superior performance over ReAct as analyzed in Section 4.2. Moreover, compared to the visual reasoning frameworks VDGD and CCoT, which also involve multiple image inputs, ProReason’s token consumption on MME is only about 20% higher, yet it achieves a 12.7% performance improvement. Regarding time efficiency, ProReason achieves 11.1% higher performance than VDGD and CCoT on MME with comparable time consumption (18.4s/sample vs. VDGD’s 12.3s and CCoT’s 12.8s), while ReAct is slower than ProReason across datasets, highlighting ProReason’s efficiency advantage. Furthermore, as discussed in Section 4.3, ProReason’s robust visual reasoning capabilities can be transferred to downstream task models, further ensuring the performance and efficiency of our approach.
ProReason achieves a balance between performance and efficiency through its Memory component and adaptive mechanisms. The Memory component enables ProReason to keep compact information representations, avoiding lengthy reasoning traces like ReAct, reducing token use and boosting efficiency. Additionally, ProReason dynamically adjusts Proactive Visual Perception iterations based on question difficulty, minimizing overhead for simple tasks while enhancing complex problem-solving. As shown in Table 12, MathVista’s greater challenge prompts more iterations compared to MME, resulting in higher token usage and longer reasoning times. This aligns with the difficulty levels of the datasets, demonstrating the adaptive nature of ProReason.
4.5 Ablation Study and Further Analysis
Relative Importance of Sub-agents.
To evaluate the importance of each sub-agent in ProReason, in Section A.5.1, we replace Dispatcher, Vision Expert, Insight Expert, Referee, and Summarizer individually with the less capable Llama3-LLaVA-NeXT-8B(which demonstrates weaker visual understanding and textual reasoning capabilities), while keeping the other sub-agents as GPT-4o-mini. The performance degradation on the MME and MMMU benchmarks is then used to measure the significance of each sub-agent. Results indicate that the Summarizer is the most critical sub-agent, closely followed by Referee.
Which One is More Crucial: Visual Understanding or Textual Reasoning
In Section A.5.2, we perform comparative experiments by substituting the vision expert and text sub-agents in the ProReason with Llama3-LLaVA-NeXT-8B and GPT-4o-mini respectively. Our findings indicate that while both visual understanding and textual reasoning capabilities are essential for multimodal tasks, textual reasoning ability holds greater significance in visual reasoning tasks. This result is consistent with our earlier analysis in Section A.5.1, which identifies the Summarizer and Referee as the most critical sub-agents.
The Critical Implication of Decoupling.
In Section A.5.3, to validate the necessity of decoupling visual perception and textual reasoning in ProReason, we systematically merge sub-agents through three configurations while preserving identical prompts and procedures, with observed performance degradation quantitatively demonstrating the critical role of decomposed processing in enhancing capabilities. The experimental results indicate that Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks
Reasoning Process Evaluation of ProReason.
In Section A.5.4, we evaluate the responses generated by ProReason using LLMs. The analysis reveals that, compared to CoT, ProReason produces more relevant answers with reduced redundancy and deficiency, consistent with its enhanced performance.
Referee’s Dispel of Hallucinations.
In Section A.5.5, adhering to the settings in Section 4.1, we assess ProReason (powered by GPT-4o-mini) on MMMU and HallusionBench with different attempt allowances (1/3/5). When attempts are unsuccessful, systematic Memory clearance is triggered (determined by the Referee’s lack of sufficient information). As the number of attempts increases, the Referee has more opportunities for information filtering. The observed performance variations illustrate the crucial influence of the Referee’s decision - making and filtering efficiency on the system’s capabilities. Experimental results indicate that the Referee module effectively filters hallucinated information to improve the visual comprehension capabilities of our framework.
Frequency of selection of various experts.
In Section A.5.6, we assess how often the Dispatcher selects the Vision Expert or Insight Expert across both MME and MMMU benchmarks. The experimental results demonstrate that ProReason adaptively adjusts the frequencies of expert selection, leading to consistent performance improvements.
5 Conclusion
In this paper, we first validate that existing multi-modal reasoning approaches still suffer insufficient and irrelevant visual descriptions, as well as limited multi-modal capacities. To address these issues, we decompose the visual reasoning process into visual perception and textual reasoning stages, and introduce a novel visual reasoning framework named ProReason, featuring decoupled vision-reasoning capabilities and multi-run proactive perception. Empirically, extensive experiments demonstrate the superiority of ProReason over both passive image information acquisition methods and multi-step reasoning frameworks for text-only LLMs across multiple visual reasoning benchmarks with both open-source and closed-source models. Notably, our method showcases the remarkable feasibility of integrating LLMs for multi-modal reasoning with dramatically improved performance, highlighting the great potential for LLM-assisted LVLM reasoning in future research.
6 Limitations
In this section, we analyze the limitations of the proposed method based on typical errors made by GPT-4o-mini-driven ProReason, as exemplified in Figures 6 and 7, to gain further understanding and identify potential research directions.
Cumulative Errors.
As illustrated in Figure 6, the vision expert mistakenly perceives the clock as 6:25, which misguides the reasoning of subsequent agents and ultimately leads to an incorrect conclusion. More broadly, similar misperceptions occur frequently in errors made by ProReason. This indicates that, with the assistance of LLMs, ProReason has effectively addressed the reasoning deficiencies in multi-modal tasks, while the vision expert plays a significant role for further improvement of multi-modal capabilities.
Contradictory Information among Agents.
Considering that multiple agents are engaged in the answering process, we try to find instances where contradictory information is provided by different agents, especially the vision expert and Insight Expert. However, as shown in Figure 6 and 7, when one agent (e.g., the vision expert) makes an error and the referee even hints at a possible mistake, other agents (e.g., the Insight Expert) tend to adhere to the available information instead of questioning it. This tendency results in a failure to find cases with contradictory information, and also highlights the importance of a reflection mechanism (Ji et al., 2023) in agent collaboration, which is left for future exploration.
We also examined existing multi-step reasoning frameworks,
such as ReAct (Yao et al., 2022), ToT (Yao et al., 2024), and Insight-V (Dong et al., 2024), aiming to find solutions for resolving accumulated errors and contradictory information. However, we found that these approaches may also fail to identify effective solutions and do not discuss the impact of the aforementioned drawbacks. We sincerely WELCOME any constructive discussions regarding accumulated errors and contradictory information! Besides, ProReason demonstrates significantly improved overall performance, suggesting fewer errors made by ProReason in the whole task distribution.
7 Ethics Statement
We adhere strictly to the ACL Code of Ethics throughout our research. To our knowledge, the methods we introduce pose no foreseeable risks. We provide comprehensive details of the computing infrastructure used for all computational experiments in the paper, along with transparent statistics on our results and a detailed configuration of our experimental setup, including the optimal hyperparameter values. Furthermore, we will release the code upon publication to facilitate easy public reproducibility.
Appendix A Appendix
A.1 Related work
Large Visual-Language Model.
Recently, large vision-language models (LVLMs) (Yang et al., 2024; Bai et al., 2023; Chen et al., 2023; Liu et al., 2024b) have garnered widespread attention and demonstrated remarkable advancements in understanding and generating multi-modal contents. In the open-source domain, numerous LVLMs, like LLaVA (Liu et al., 2023c, b, 2024a; Li et al., 2024a, b) and InternVL (Chen et al., 2024c) families, have been extensively developed. In the closed-source domain, proprietary models such as GPT-4o [^5] https://openai.com/index/hello-gpt-4o/ and Gemini Pro 2.5666https://deepmind.google/technologies/gemini/pro/ have also achieved significantly success. Additionally, multi-agent frameworks like VipAct (Zhang et al., 2024c) have been developed to improve LVLMs’ perception of visual details. Despite these advancements, existing LVLMs still encounter challenges in effectively integrating visual understanding with textual reasoning capabilities simultaneously. This limitation is particularly evident in their diminished attention to image content during visual reasoning process, such as chart interpretation and visual math reasoning, leading to degraded performance (Liu et al., 2023a; Ghosh et al., 2024) and motivating more effective solutions.
Multi-step Reasoning Framework.
Multi-step reasoning frameworks improve LLM performance by breaking down complex tasks. Chain-of-Thought (CoT) (Wei et al., 2022) enhances reasoning via explicit intermediate steps, demonstrating effectiveness in both textual and visual tasks (Zhang et al., 2024b), while Tree-of-Thoughts (ToT) (Yao et al., 2024) extends this by evaluating multiple reasoning paths. ReAct (Yao et al., 2022) integrates dynamic knowledge retrieval during reasoning. In visual reasoning, several frameworks assist LVLM by extracting image information into text. Visual Description Grounded Decoding (VDGD) (Ghosh et al., 2024) describes the image and appends this description to the prompt, aiding LVLMs in visual reasoning tasks. Compositional Chain-of-Thought (CCoT) (Mitra et al., 2024) guides LVLMs to create scene graphs (SGs) that link visual and textual domains, supporting subsequent tasks. However, these methods often use a question-agnostic, reasoning-free visual extraction process, resulting in irrelevant or redundant information. Insight-V (Dong et al., 2024) trains LVLM with a multi-agent system, but all agents rely on a single LVLM, blending visual understanding with textual reasoning and failing to address the challenge of effectively managing both. Unfortunately, this method does not release its code and prompts, resulting in low reproducibility and making comparison difficult. In response to these drawbacks, we introduce ProReason, which decouples visual reasoning tasks into proactive visual perception (i.e., eyesight) and textual reasoning (i.e., wisdom), and makes all prompts available. By leveraging the strengths of the decoupled system, ProReason effectively integrates existing powerful LLMs to achieve high-performance visual reasoning and successfully transfers this capability to downstream task models.
A.2 Dataset
To thoroughly assess the performance of our framework, we have carried out experimental evaluations using four benchmark datasets: the Multi-modal Large Language Model Evaluation (MME) (Yin et al., 2023), the cross-disciplinary Massive Multi-modal Understanding and Reasoning benchmark (MMMU) (Yue et al., 2023), the visual mathematical reasoning assessment MathVista (Wang et al., 2024), and the multimodal illusion detection benchmark HallusionBench (Liu et al., 2023a). Each of these benchmarks necessitates strong visual reasoning capabilities for successful task completion, and we provide concise descriptions below:
  • MME is an inclusive benchmark that encompasses 14 subtasks, designed to evaluate perceptual and cognitive abilities. Given our focus on visual reasoning, we select the cognition-relevant tasks, including Commonsense Reasoning, Numerical Calculation, Text Translation, and Code Reasoning.
  • MMMU evaluates multi-modal models with multidisciplinary tasks that require college-level domain-specific knowledge and detailed reasoning. It comprises 11,500 questions across 30 disciplines and 183 sub-fields, emphasizing advanced perception and domain-specific reasoning.
  • MathVista focuses on more challenging mathematical reasoning tasks that demand precise visual recognition and compositional reasoning. It includes 6,141 examples from 31 multi-modal mathematics datasets.
  • HallusionBench evaluates models’ ability to reason with images such as statistical charts, emphasizing nuanced visual understanding. It consists of 346 images paired with 1,129 questions, meticulously crafted by experts.
A.3 Baselines
  1. Direct. As indicated by the name, models are required to answer questions directly without dedicated prompts. This baseline is set to evaluate the initial performance of base models.
  2. CoT. CoT is an advanced prompting method that encourages LLMs to break complex tasks down into a series of easy steps, which has been applied broadly and verified to boost the reasoning performance remarkably (Chu et al., 2023).
  3. ReAct. ReAct is an LLM-specific agent framework, which performs tasks by alternating between reasoning and execution behaviors. To extend it to multi-modal domain, we use two LVLMs to perform both steps, and rename them as the Vision and Insight Experts, respectively. This aligns with our notions for easy understanding, and is shown in Figure 1.
  4. VDGD. VDGD involves two main steps: initially, LVLMs generate detailed image captions, which are then incorporated into prompts to aid inference. During the inference process, VDGD also utilizes a formula based on Kullback-Leibler divergence to select tokens that minimally deviate from the description, thereby enhancing the relevance of the model’s reasoning to the image [^7] Since we cannot obtain the tokens output by GPT-4o-mini, we omit the step of selecting the token with the smallest deviation from the image description when implementing VDGD for GPT-4o-mini..
  5. CCoT. Given an image and the question, CCOT first generates a scene graph of the image with LVLMs, and then extracts the answer by prompting the LVLMs with the graph.
  6. R1-Onevision. R1-Onevision-7B is an advanced multimodal reasoning model based on Qwen2.5-VL-7B-Instruct that has deep thinking capabilities akin to Deepseek-R1 by transforming images into structured textual representations and employing a training framework that merges supervised fine-tuning with reinforcement learning.
A.4 Model Training
[…]
A.5 Supplementary Results and Analysis
Table 4: Performance of three recent LVLMs on MMMU dataset with different assisting techniques.
Model Method
Direct CoT VDGD
Llama3-LLaVA-NeXT-8B 41.8 41.5 42.7
Qwen2.5-VL-7B-Instruct 51.8 52.2 52.6
Table 5: Effectiveness evaluation of passive captions along Detail Level, Question Relevance, and Reasoning Effective Info Inclusion. “True” and “False” denote the response correctness of Llama3-LLaVA-NeXT-8B.
Score Llama3-LLaVA-NeXT-8B
True False
Detail Level 4.43 3.93
Question Relevance 3.87 3.30
Reasoning Effective Info Inclusion 3.91 3.57
A.5.1 Relative Importance of Sub-agents
To assess the importance of each sub-agent within the ProReason framework for visual reasoning tasks, we design five scenarios where Llama3-LLaVA-NeXT-8B acts as Dispatcher, Vision Expert, Insight Expert, Referee, or Summarizer, respectively, while the other sub-agents are powered by GPT-4o-mini. Given that Llama3-LaVA-NeXT-8B exhibits weaker visual understanding and textual reasoning capabilities than GPT-4o-mini, the more significant the performance drop incurred by replacing a sub-agent with Llama3-LaVA-NeXT-8B is, the more important that sub-agent is. Here we primarily consider the MME and MMMU benchmarks due to their comprehensive question coverage. The experimental results are presented in Table 6.
Summarizer is the most crucial sub-agent, closely followed by Referee. The replacement of Summarizer results in the most notable performance decline on both MME and MMMU tasks, reaching 6.2% and 10.6%, respectively. This highlights the critical function of the Summarizer in integrating all available information to conclude final answers. Besides, the substitution of Referee leads to a 10.1% reduction on MMMU. Given that MMMU is more challenging than MME, this finding underscores the essential role of the Referee in assessing the sufficiency of information, particularly in more complex visual reasoning tasks. The analysis in Section A.5.5 also demonstrates that the Referee plays a crucial role in enabling ProReason to accurately interpret visual detail information.
Relatively, Dispatcher and Insight Expert are the least essential sub-agents. Specifically, despite a decline, these two sub-agents exhibit significantly less performance degradation than other sub-agents. This can be attributed to the easier task of the Dispatcher, which requires minimal textual reasoning capabilities, and the infrequent calls of the Insight Expert, which is only activated when additional information needs to be inferred—a situation that is rare in current benchmarks. Besides, both sub-agents operate within the acquisition loop, allowing for greater error tolerance. Even if some error occurs, subsequent iterations can compensate for the missing information.
In summary, each sub-agent contributes to the performance of ProReason, underscoring their necessity. Relatively, the Summarizer and Referee are the most critical sub-agents, while the Dispatcher and Insight Expert have the least impact.
Table 6: Performance of ProReason across five scenarios for sub-agent assessment on visual reasoning tasks. For each scenario, one sub-agent is replacing with Llama3-LLaVA-NeXT-8B, while the others are performed by GPT-4o-mini. The blue text indicates the performance decline compared to the scenario with all agents performed by GPT-4o-mini.
Model Agent Dataset
MME MMMU
GPT-4o-mini 90.4 61.6
Llama3- LLaVA- NeXT-8B Dispatcher 88.8 (-1.6) 60.9 (-0.7)
Vision Expert 84.7 (-5.7) 54.5 (-7.1)
Insight Expert 88.7 (-1.7) 60.2 (-1.4)
Referee 89.6 (-1.1) 51.5 (-10.1)
Summarizer 84.2 (-6.2) 51.0 (-10.6)
A.5.2 Which One is More Crucial: Visual Understanding or Textual Reasoning?
ProReason effectively decouples the visual understanding and textual reasoning capabilities of LVLMs. However, it remains unclear which of these two capacities is more critical for visual reasoning tasks. To answer this question, we conduct comparative experiments of the following three scenarios:
  1. Llama3-LLaVA-NeXT-8B as All Sub-Agents. All sub-agents within ProReason framework are performed by Llama3-LLaVA-NeXT-8B model.
  2. GPT-4o-mini as Vision Expert. Based on the above scenario, we implement the Vision Expert with GPT-4o-mini, while keep the other textual sub-agents unchanged.
  3. GPT-4o-mini as Textual Sub-Agents. Reversely, we utilize Llama3-LLaVA-NeXT- 8B as the Vision Expert, and GPT-4o-mini for the other vision-irrelevant sub-agents.
Textual reasoning capabilities outweigh visual understanding for multi-modal reasoning tasks, although both are important. As shown in Table 7, replacing either the Vision Expert or the other agents with the more capable GPT-4o-mini achieves consistent performance enhancement, highlighting the significance of both capabilities. However, substituting the textual sub-agents with GPT-4o-mini results in a more substantial performance boost compared to replacing the Vision Expert. This underscores the greater importance of textual reasoning over visual understanding for multimodal reasoning tasks, aligning with our previous analysis in Section A.5.1 that identifies the Summarizer and Referee as the most crucial sub-agents.
Table 7: Performance of ProReason with different configurations for the relative importance assessment between visual understanding and textual reasoning capabilities on visual reasoning tasks. The red text highlights the performance improvements brought about by the introduction of GPT-4o-mini.
Model Agent Dataset
MME MMMU
GPT-4o -mini Textual Sub-Agents 84.7 (+13.2) 54.5 (+2.0)
Vision Expert 77.8 (+6.3) 53.4 (+0.9)
Llama3- LLaVA- NeXT-8B All Sub-Agents 71.5 52.5
COT 58.8 41.5
Direct 61.5 41.8
A.5.3 The Critical Implication of Decoupling
To validate the critical implication of decoupling visual perception and textual reasoning in ProReason while eliminating potential confounding factors from prompt engineering and multiple CoT implementations, we systematically integrate sub-agents through three configurations:
  1. Merge the Vision Expert and Insight Expert into a single sub-agent to examine the necessity of modality decoupling during sub-task execution.
  2. Integrate the Dispatcher, Vision Expert, Insight Expert, and Referee as a unified agent to verify the essentiality of the process design for the Proactive Visual Perception stage.
  3. Fully consolidate all five original sub-agents to demonstrate the pivotal role of decomposing visual reasoning tasks into distinct Proactive Visual Perception and Textual Reasoning phases.
The merged agents preserve identical prompts and maintain the same task execution procedures as their original counterparts, with their prompts shown in Figures 13, 14, and 15. Therefore, the performance degradation caused by agent merging quantitatively demonstrates the critical implication of decoupled processing in enhancing ProReason’s capabilities.
Decoupling serves as a crucial mechanism for improving ProReason’s performance in complex visual reasoning tasks. As shown in Table 8, merging Vision and Insight Experts results in a 3% performance drop on the MMMU benchmark, while combining Proactive Visual Perception with Textual Reasoning leads to a more significant 4.8% reduction (56.8 vs. CoT’s 58.5). Although agent merging also causes performance declines on MME, the merged versions still outperform CoT. Notably, given MMMU’s substantially higher complexity compared to MME, these findings reveal that the decoupling of visual-textual processing fundamentally drives its performance gains in complex scenarios.
Table 8: Impact of decoupling visual perception and textual reasoning on ProReason performance with results shown from sub-agent consolidation experiments. The blue text highlights the performance degradation due to sub-agent integration.
GPT-4o-mini Dataset
MMMU MME
ProReason 61.6 91.9
CoT 58.5 87.8
Vision & Reasoning
Expert Integration
58.6 (-3.0) 90.4 (-1.5)
Dispatcher,
Vision & Insight Expert,
& Referee Integration
57.8 (-3.8) 90.8 (-1.1)
All Sub-Agents Integration 56.8 (-4.8) 89.2 (-2.7)
A.5.4 Reasoning Process Evaluation of ProReason
Inspired by Liu et al. (2023c), we design a pipeline using LLMs to further analyze the performance of ProReason. Given that GPT-4o achieved a 70.3% score on MMMU, significantly surpassing GPT-4o-mini’s 58.5% and ProReason driven by GPT-4o-mini’s 61.6%, we adopt GPT-4o’s CoT-based answers on MMMU as the standard answer. Using this standard answer, we evaluate the reasoning process of GPT-4o-mini (CoT), the Memory of ProReason powered by GPT-4o-mini (ProReason-Memory), and the reasoning process of ProReason’s Summarizer (ProReason-Summarizer) on MMMU. This assessment focuses on three key metrics: the relevance to standard answer (RE ↑), the degree of redundant information (RI ↓), and the extent of missing information (MI↓), where arrows indicate the directions of improvement. The evaluation process is driven by GPT-4o and the prompt is shown in Figure 12.
Specifically, compared to GPT-4o-mini (CoT), ProReason-Summarizer produces more relevant answers with less redundancy and deficiency, aligning with its improved performance. Compared to ProReason-Summarizer, ProReason-Memory exhibits the same RE, higher RI and lower MI scores. This suggests that ProReason allows some redundancy to prevent information loss in memory, as the former typically leads to more serious consequences than the latter. Subsequently, Summarizer can leverage its powerful reasoning capabilities to select the most relevant memory.
Table 9: Performance of ProReason driven by GPT-4o-mini assessed by LLMs compared to CoT on MMMU benchmark. Mainly includes three key metrics: the relevance to standard answers (RE ↑), the degree of redundant information (RI ↓), and the extent of missing information (MI ↓), where arrows indicate the directions of improvement.
GPT-4o-mini Metrics
RE ↑ RI ↓ MI ↓
COT 4.67 3.33 1.40
ProReason-Memory 4.83 3.66 1.17
ProReason-Summarizer 4.83 2.88 1.33
A.5.5 Referee’s Dispel of Hallucinations
Following the implementation details outlined in Section 4.1, we evaluate the ProReason system powered by GPT-4o-mini on both MMMU and HallusionBench under varying attempt allowances (i.e., 1, 3, and 5 attempts). Every unsuccessful attempt reflects the Referee’s persistent determination that the information stored in Memory is insufficient to solve the problem. Before each new attempt, Memory is systematically cleared, ensuring the removal of information deemed irrelevant by the Referee’s assessment. Consequently, increased attempt allowances essentially empower the Referee with enhanced opportunities for information filtration. The observed performance variations of ProReason across different attempt quotas demonstrate the critical impact of Referee’s decision-making mechanism and information filtering efficacy on system capability.
The Referee module effectively filters hallucinated information to enhance the visual comprehension capabilities of our framework. As demonstrated in Table 10, the performance improvement on HallusionBench (2.8%) significantly outpaces that on MMMU (1.5%) as attempt opportunities increase from 1 to 5. Given HallusionBench’s dual emphasis on reasoning proficiency and precise evaluation of visual details/hallucination control compared to MMMU, these results suggest that the Referee mechanism can effectively identify erroneous or irrelevant visual information, thereby strengthening ProReason’s capacity for meticulous visual understanding.
Table 10: Impact on HallusionBench and MMMU performance across different attempt allowances.
Dataset Attempt Allowances
1 3 5
MMMU 60.1 59.9 61.6
HallusionBench 57.1 58.9 59.9
A.5.6 Frequency of selection of various experts
As listed in Table 11, we evaluate the frequency of the Dispatcher choosing the Vision Expert or Insight Expert on both MME and MMMU benchmarks, with MMMU requiring higher visual and reasoning abilities. Specifically, compared to MME, the frequencies for both the Vision and Insight Experts are higher on the MMMU benchmark, aligning with their difficulty levels. Together with the results in Table 1 of our submission, ProReason can adaptatively increase the frequencies of experts, and provide consistent performance improvements (i.e., 11.2% and 13.2%). Despite the lower frequency of the Insight Expert, the significant performance enhancement highlights the importance of LLM-assisted reasoning capabilities for reasoning-essential questions. Additionally, the frequency of the Vision expert exceeding 1 underscores the importance of referees, which controls the loop to call experts multiple times, alleviating the issue of insufficient information.
Table 11: Frequency of the Dispatcher choosing the Vision Expert or Insight Expert on both MME and MMMU benchmarks.
Dataset GPT-4o-mini
Vision Expert Insight Expert
MME 1.16 0.12
MMMU 1.64 0.38
A.5.7 Performance Evaluation of ProReason on Supplementary Datasets
To further assess the effectiveness of ProReason across diverse visual tasks, we expand upon the four datasets introduced in Section 4.1 by incorporating three additional benchmarks: MathVerse (Zhang et al., 2024a), MMStar (Chen et al., 2024a), and A-OKVQA (Schwenk et al., 2022). MathVerse focuses on visual reasoning within mathematical contexts, while MMStar highlights visual dependency and data reliability. A-OKVQA, in contrast, serves as a knowledge-based VQA dataset requiring minimal reasoning. For MathVerse, we utilize the Text Lite subset, which prioritizes visual reasoning by minimizing textual content. In MMStar, subsets related to coarse and fine-grained perception that are unrelated to reasoning are excluded. Table 13 presents ProReason’s performance on each dataset, showcasing significant improvements across various forms of multi-modal tasks and highlighting the effectiveness and generalizability of the perception and reasoning decoupling approach.
Table 12: Number of iterations required for each problem during the Proactive Visual Perception phase on MME and MathVista benchmarks.
Model Dataset
MME MathVista
GPT-4o-mini 1.28 2.13
Table 13: Performance of multiple approaches with GPT-4o-mini across 7 visual benchmarks. “Hallu.” is the abbreviation of HallusionBench. Based on the performance of the direct method, red and blue signify the improvement and degradation, respectively.
Model Method Dataset
MME MMMU MathVista Hallu. MathVerse MMStar A-OKVQA
GPT-4o-mini Direct 79.2 48.4 53.0 56.0 28.2 46.2 78.6
VDGD 82.3 (+3.1) 51.4 (+3.0) 51.2 (-1.8) 52.4 (-3.6) 30.1 (+1.9) 47.0 (+0.8) 79.4 (+0.8)
CCoT 80.8 (+1.6) 54.2 (+5.8) 53.6 (+0.6) 56.7 (+0.7) 29.2 (+1.0) 45.4 (-0.8) 79.2 (+0.6)
CoT 87.8 (+8.6) 58.5 (+10.1) 53.8 (+0.8) 56.3 (+0.3) 28.9 (+0.7) 47.2 (+1.0) 80.9 (+2.3)
ReAct 87.3 (+8.1) 54.8 (+6.4) 49.3 (-3.7) 51.1 (-4.9) 30.4 (+2.2) 46.7 (+0.5) 80.6 (+2.0)
ProReason 91.9 (+12.7) 61.6 (+13.2) 54.9 (+1.9) 59.9 (+3.9) 31.6 (+3.4) 49.1 (+2.9) 81.3 (+2.7)
Average 84.63 54.82 52.63 55.4 29.7 46.9 80.0
Table 14: Different model configurations of Textual Sub-Agents for implementing LLM-assisted visual reasoning in ProReason. The abbreviations 4o-mini, Qwen72B, and Qwen3 refer to GPT-4o-mini, Qwen2.5-72B-Instruct, and Qwen3-32B, respectively. "Assisted" stands for LLM-assisted reasoning.
Model Method Vision Expert Summarizer Other Textual Sub-Agents
Llama3-LLaVA-NeXT-8B ProReason + 4o-mini Assisted Llama3-LLaVA-NeXT-8B GPT-4o-mini GPT-4o-mini
ProReason + Qwen72B Assisted Llama3-LLaVA-NeXT-8B Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct
Qwen2.5-VL-7B-Instruct ProReason + Qwen72B Assisted Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct
ProReason + Qwen3 Assisted Qwen2.5-72B-Instruct Qwen3-32B Qwen2.5-72B-Instruct
A.6 Demonstrative Examples
Figure 2: An example with three reasoning frameworks: fine-grained caption, chain-of-thought, and ProReason. ProReason enables LVLMs to proactively acquire necessary information in a question-oriented manner, and predicts answers based on the collected information. Apparently, ProReason is superior to previous methods, which often describe question-irrelevant visual details, or overlook informative elements. Green indicates correct information or conclusions, while red signifies incorrect ones.
Figure 3: A complete reasoning process of ProReason for the case shown in Figure 2.
Figure 4: Additional examples of images-unrelated Chain-of-Thought reasoning.
Figure 5: Additional examples of question-unrelated passive image captions.
Figure 6: A typical mistake made by ProReason. The vision expert incorrectly identifies 4:30 as 6:25, leading the other agents to base their judgments on this erroneous information, and ultimately resulting in the wrong conclusion.
Figure 7: A typical mistake made by ProReason. The vision expert incorrectly provided the coordinates of the three vertices of the triangle, leading the other agents to base their judgments on this erroneous information, and ultimately resulting in the wrong conclusion.
A.7 Prompts
Figure 8: Prompt templates of Relevance Score and caption effectiveness evaluation.
Figure 9: Prompt templates of Chain-of-Thought, fine-grained image captions generation, Compositional Chain-of-Thought (CCoT), and ReAct.
Figure 10: Prompt templates of Dispatcher.
Figure 11: Prompt templates of Vision Expert, Insight Expert, Referee, and Summarizer.
Figure 12: Prompt templates of LLMs-based ProReason evaluation.
Figure 13: Prompt templates of Vision and Insight Expert Integration.
Figure 14: Prompt templates of Dispatcher, Vision Expert Insight Expert and Referee Integration.
Figure 15: Prompt templates of All Five Sub-Agent Integration.